### v1

Abstract
Many incurable diseases prevalent across global societies stem from various influences, including lifestyle choices, economic conditions, social factors, and genetics. Research predominantly focuses on these diseases due to their widespread nature, aiming to decrease mortality, enhance treatment options, and improve healthcare standards. Among these, kidney disease stands out as a particularly severe condition affecting men and women worldwide. Nonetheless, there is a pressing need for continued research into innovative, early diagnostic methods to develop more effective treatments for such diseases. Recently, automatic diagnosis of Kidney Cancer has become an important challenge especially when using deep learning (DL) due to the importance of training medical datasets, which in most cases are difficult and expensive to obtain. Furthermore, in most cases, algorithms require data from the same domain and a powerful computer with efficient storage capacity. To overcome this issue, a new type of learning known as transfer learning (TL) has been proposed that can produce impressive results based on other different pre-trained data. This paper presents, to the best of the authors’ knowledge, the first comprehensive survey of DL-based TL frameworks for kidney cancer diagnosis. This is a strong contribution to help researchers understand the current challenges and perspectives of this topic. Hence, the main limitations and advantages of each framework are identified and detailed critical analyses are provided. Looking ahead, the article identifies promising directions for future research. Moving on, the discussion is concluded by reflecting on the pivotal role of TL in the development of precision medicine and its effects on clinical practice and research in oncology.
Index Terms:
Artificial intelligence, Deep learning, Domain adaptation, Fine-tuning, Kidney cancer, Transfer learning.
[…]
=-15pt
I Introduction
The kidneys play a crucial role in regulating the body’s fluid balance by filtering out toxins and excreting waste products. Various disorders, such as kidney cancer (KC), can impair kidney function [1]. Fig.1 shows the estimated number of KC incidence and mortality in 2022, for both genders in the world [2]. KC is typically graded on a scale of 1 to 4, with 1 being the least aggressive and 4 being the most aggressive. The grade of a KC is determined by analyzing the cancer cells under a microscope and assessing the degree of differentiation and anaplasia. Grade 1 known as papillary renal cell carcinoma (PRCC), consists of well-differentiated cells closely resembling normal kidney cells [3]. These cancers tend to grow slowly. Grade 2 or chromophobe renal cell carcinoma (CRCC), comprises moderately differentiated cells with some resemblance to normal kidney cells. These cancers grow more slowly than grade 3 or 4 cancers. Grade 3 known as clear cell renal cell carcinoma (CCRCC), consists of poorly differentiated cells bearing little resemblance to normal kidney cells [4]. These cancers tend to grow and spread more rapidly than grade 1 or 2 cancers. Grade 4 called as sarcomatoid renal cell carcinoma (SRCC), is a rare and highly aggressive type of KC that grows and spreads rapidly. It’s important to note that the grade of a KC is one of several factors considered when determining the best course of treatment [5, 6].
Figure 1: The estimated number of KC incidence and mortality in the world.
AL
artificial learning
AI
artificial intelligence
CNN
convolutional neural network
CT
computerized tomography
CRCC
chromophobe renal cell carcinoma
CCRCC
clear cell renal cell carcinoma
CAD
computer-aided detection
DA
domain adaptation
DL
deep learning
DNN
Deep neural network
F1
f1-score
HKS
human kidney segmentation
KC
kidney cancer
KiTS19
kidney tumor segmentation 19
LLMs
Large language models
ML
machine learning
MRI
magnetic resonance imaging
MPS
main processing steps
PRCC
papillary renal cell carcinoma
RCC
renal cell carcinomas
SVM
support vector machine
SRCC
sarcomatoid renal cell carcinoma
TL
transfer learning
TCGA
the cancer genome atlas
US
ultrasound
UTL
unsupervised TL
CTC
connectionist temporal classification
BERT
bidirectional encoder representations from Transformer
XAI
explainable AI
Moreover, KC can be identified by employing a range of imaging methods such as magnetic resonance imaging (MRI), X-rays, and ultrasound (US) imaging [7]. These methods are capable of generating comprehensive visuals of the kidneys, enabling physicians to spot any irregularities or tumors [8]. Nonetheless, despite the utility of medical imagery in disease diagnosis, its reliability is not absolute in certain scenarios. Particularly in cancer cases, more precise techniques may be required to aid physicians in detecting malignant cells and assessing their metastasis to other body regions [9]. Moreover, interpreting whether an image truly signifies the presence of cancer can be challenging. Medical imagery is often rich in detail and data, complicating its analysis. In addition, the overwhelming volume of images and visual assessments may contribute to medical inaccuracies. The burden of analyzing numerous images can diminish a radiologist’s focus and their capacity to notice crucial nuances. Such circumstances can lead to misdiagnosis, diagnosis delays, or even inappropriate treatments [9].
In recent years, deep learning (DL)-based early detection methods have recorded significant achievements in identifying various terminal illnesses [10, 11, 12]. These methods focus on accurately extracting diverse features that assist medical professionals in the diagnostic process through the utilization of the network’s multiple layers [13, 14]. Attaining the necessary diagnostic precision for incurable diseases affecting the kidneys presents challenges due to: (i) the insufficiency of data resources for the training phase [15, 16]; (ii) the time-consuming, expensive, and exhaustive nature of data collection [17, 18]; (iii) the requirement for data annotation in the training phase; and (iv) the impracticality of manually labeling large datasets by experts, which is often unfeasible [19, 20].
While many DL algorithms, such as the convolutional neural network (CNN) [21, 11, 13, 10, 22, 23], are proficient at recognizing complex patterns, they encounter significant challenges, particularly during the transition from training to testing phases, which can impede their effectiveness. Additionally, various studies [24] have pointed out that the majority of datasets utilized for diagnosing either early or advanced cancer stages are often inadequate for the diagnostic process. This inadequacy is due to the intricate structure of certain organs, like the kidneys, and the variations in image depth, making it exceedingly difficult to accurately identify the type or grade of cancer [25, 26, 27]. Furthermore, these algorithms heavily rely on the assumption that the training and testing data are extracted from identical feature spaces and distributions. A deviation between these can necessitate a comprehensive model reconfiguration, which is not only costly but also labor-intensive. As [28] notes, supervised learning algorithms can yield remarkable outcomes with extensive volumes of annotated data. Yet, they generally falter with the complex variations present in medical images, where labeled data are scarce, and backgrounds are complex. Conversely, diagnostic systems powered by artificial intelligence (AI) critically depend on having access to well-distributed and annotated datasets for training, a requirement that is often hard to fulfill [29]. This underscores the need for enhancements in AI-based diagnostic systems for incurable diseases, aiming to simplify them to improve diagnostic efficiency, timing, and precision [30].
To meet these goals, efforts are being directed towards leveraging transfer learning (TL) in many research fields [31, 32, 33, 34, 35]. This process fundamentally involves repurposing a network that has been previously trained on a specific dataset for a distinct task. This entails employing established models such as LeNet, AlexNet, VGGNet, ResNet, GoogLeNet, DenseNet, XceptionNet, and SqueezeNet [32] for tasks on smaller datasets, thereby facilitating the omission of labels. This concept draws inspiration from the human ability to apply knowledge and experiences from past learning to foster new insights. Consequently, various databases hosting extensive image collections have been made available online, such as the CNN algorithm’s application on the ImageNet dataset, which is accessible for public experimentation with DL-based algorithms. This approach effectively mitigates data scarcity issues and circumvents the need for extensive data annotation [36]. Moreover, this algorithmic approach alleviates the lengthy durations typically required for learning processes, which can extend over days or weeks. In the realm of diagnosing terminal illnesses, TL has been categorized into two methodologies: fine-tuning and feature extraction, both of which have shown promising outcomes, notably in KC detection. Numerous studies within this domain, for instance, [37], exemplify its application. TL is particularly beneficial for cancer detection and diagnosis, including KC [38, 39]. Utilizing a pre-trained DL model, like a CNN, fine-tuned on kidney imaging scans (e.g., computerized tomography (CT) or MRI), can enhance its efficacy in identifying and diagnosing kidney tumors (Fig. 2). A method for fine-tuning a model for KC detection might involve using a dataset comprising both healthy and cancerous kidney scans, adjusting the model’s final layers’ weights to better recognize cancerous areas. Additionally, the model could be optimized by either freezing specific layers while training the last few or starting anew with the pre-trained weights. An alternative TL strategy in KC detection could leverage a model pre-tuned for a similar task, such as lung cancer detection, and further fine-tune it on kidney scans, beneficial when large datasets for KC detection are unavailable. TL proves invaluable in medical imaging, significantly when large, labeled datasets are scarce, enhancing the efficiency and accuracy of kidney tumor identification and diagnosis.
Figure 2: Kidney classification based on TL.
I-A Our Contributions
Numerous reviews have covered the contributions of the TL algorithm, citing works such as [37, 40, 41]. However, this paper offers a novel examination by showcasing the role of TL in identifying and managing various types and degrees of renal conditions within a unified and comprehensive framework that encompasses classification and concepts. This approach aids in demystifying the domain of TL and its utilization in KC detection endeavors. The primary focus has been on the enhancement through fine-tuning and domain adaptation (DA) in KC detection, particularly highlighting recent cross-domain advancements. Furthermore, this review underscores the significance of employing TL in data fusion for improving cancer detection and outlines the challenges and future research avenues, suggesting a shift towards generalized TL models for KC detection. The unique contributions of this analysis include:
  • Exploring various facets of TL and DA that have propelled their evolution.
  • Offering a detailed categorization of TL and DA frameworks along with an exploration of associated concepts.
  • Conducting a thorough literature review on TL and DA’s role in KC detection, grounded in specific standards.
  • Illustrating how TL and DA frameworks enhance integrated data approaches to KC detection.
  • Highlighting unresolved issues in cancer detection faced by TL and DA, such as (1) model precision concerns, (2) selection of evaluation metrics, (3) the challenge of transferring knowledge irrelevant to the subject matter, etc.
  • Proposing future research trajectories for TL and DA frameworks that promise lower computational demands and expenses.
  • Marking this review as a pioneering effort in discussing TL and DA for KC detection, it distinctively contributes by (1) delving into the diverse applications of TL and DA, (2) integrating TL and DA within the KC detection landscape, (3) identifying assorted challenges, and (4) suggesting future research directions.
Table. I, displays the results of the comparison of the contribution of the proposed review with other surveys in the field of TL.
TABLE I: A summary of a comparison of the proposed study’s contribution with other studies in the field of TL is shown. A check mark (✓) signifies that the area has been covered, while a cross mark () indicates that it has not been addressed.
Review Year Description TLB TLS DA TLA TLD TLAD TLL KCH FD
[42] 2018 TL for mental health
[43] 2018 TL for cancer detection
[44] 2019 TL for breast cancer
[45] 2020 TL for medical image
[46] 2021 TL for EEG signal analysis
[47] 2021 TL for MRI brain
[48] 2021 TL for Alzheimer’s disease
[49] 2022 TL for neuroimaging analysis
[50] 2022 TL for KC
[51] 2023 TL for COVID-19 medical image
[52] 2023 TL for Diabetic Retinopathy
Our 2024 TL for KC
Abbreviations: TL background (TLB), TL segmentation (TLS), TL applications (TLA), TL datasets (TLD), TL advantages (TLAD), TL limitations (TLL), Key challenges (KCH), Futur direction (FD)
I-B Bibliometric Analysis
To explore and scrutinize the scientific investigations detailed in this review, bibliometric methods were utilized. The cumulative interest in TL-based KC over time is summarized in Fig. 3. This illustration provides a snapshot of the enduring interest in research on TL-based KC, highlighting a notable uptick in enthusiasm for crafting TL-based KC methodologies. Specifically, this surge in interest is represented in Fig. 3(a), indicating a sharp rise in the publication count, which peaked at 66 articles in 2023. Furthermore, Fig.3(b) displays the countries contributing most significantly to research outputs in this area, with China and the United States emerging as the most prolific in TTL-based KC research advancements. Furthermore, Fig. 3(c) identifies the differents domain of application of KC-oriented TL. Lastly, Fig. 3(d) delineates the distribution of paper types, revealing that journal articles predominate, constituting 65.8% of all published work, with conference papers making up 19.7%.
Figure 3: Bibliometric analysis in terms of (a) The number of published articles, (b) The most country known by active researchers, (c) The percentage of the contribution in each domain, and (d) Distribution of paper types.
Table. II, presents a summary of TL research conducted in the field of kidney disease diagnosis.
TABLE II: Summary of TL research conducted in the diagnosis of kidney disease.
Ref. Year Dataset Sample/Input size Used method Performance assessment Objects Methods description KSO
[53] 2022 KiTS - 3D-MS ACC:0.9390 Improve segmentation in Multi-scale approach with a Deep neural network (DNN) Kidney
-RFCNN kidney tumors. based on 3D-MS-RFCNN is applied to capture global
contextual features for kidney tumors
[54] 2022 Private 12,446 VGG16 ACC:99.30 Perform efficient an Development an intelligent system based on EANet, CCT, Kidney
data auto-diagnose kidney diseases Swin transformers and Resnet, VGG16, and Inceptionv3 models
to auto-diagnose kidney stones, cysts, and tumors.
[55] 2022 Private 158 U-Net+TL ${DC}:89.34\pm 5.31$ Improve kidney segmentation Proposed a fully automated 2D attention U-Net model for kidney Kidney
data segmentation on T1 weighted-nephrographic phase contrast enhanced
[56] 2022 Private 162 images CNN+CM DC: 0.853 segmentation of kidney Propose a cross-modal TL from CT to US to Kidney
data +TL in US images segment kidney organ based on CycleGAN, U-Net, U-Res
, PSPNet, and DeepLab v3+ and CNN models.
[57] 2021 Private 258 WSI CNN+TL SEN:0.93 Quantify non-sclerotic and A 9-layer CNN based on the common U-Net Kidney
data F1:0.87 sclerotic glomeruli from kidney. architecture was developed and tested for the segmentation
of non-sclerotic and sclerotic glomeruli of kidney.
[58] 2021 Private - VGG16+TL F1:97.03 Extract the kidney stone The U-Net was used to extract the kidney stone area. Kidney
data fragment area The VGG16 is used as the encoder of U-Net model
to extract the semantic information.
[59] 2020 Private 560/ VGG-16 ACC:0.89 Segmentation of kidneys Investigates the performances of pre-trained U-Net model Kidney
data 768X512 SEN:0.81 from 2D US images. using various backbones for segmentation of
SPE:0.92 kidneys from 2D US images.
[60] 2020 Private 196 FCNN DC:0.88 Segment organs from Used 3D U-Net kernels for segmenting 3D MRI. Kidney
data $VD:0.09\pm 0.05$ large biomedical 3D images
[61] 2019 CHOP 185 kidney TL+CNN ${ACC}:0.989\pm 0.006$ Segment the kidneys Used a boundary distance regression network Kidney
images +BDRN ${DC}:0.94\pm 0.03$ to learn kidney boundary
$MD:2.72\pm 1.61$
[62] 2019 Private 50:CAKUT TL+SVM ACC: 87.00 Classify of normal children support vector machine (SVM) classifiers and TL are used to Kidney
data 50:Controls SEN:86.00 and those with CAKUT. classify diseased and normal kidneys.
[63] 2019 KiTS19 - nnU-Net DC: 0.9620 Kidney segmentation nnU-Net algorithm used to for kidney segmentation Kidney
in CT scans
[64] 2018 Private 48 Slide CNN F1:0.8475 Classified non-sclerosed CNN model is used to classified classifies normal Glomeruli
data images and sclerosed glomeruli. and sclerotic glomeruli
[65] 2018 Private 512x512 Kid-Net F1:Artery:0.72 Segment and classify Propose Kid-Net model for semantic Kidney
data F1:Vein:0.67 kidney vessels kidney segmentation.
[66] 2018 GCDB 935 images DNN+TL ACC:0.8823 Classifying categories of Test DNN+TL model for classifying Glomeruli
SEN:0.8823 glomeruli normal and abnormal categories of glomeruli
[67] 2018 CHOP 50:CAKUT CNN+BDRN $MD:2.8717\pm 2.0665$ Segment the kidneys in Propose subsequent boundary distance regression Kidney
50:CAKUT +PWC ${SEN}:0.9354\pm 0.0625$ clinical US images to segment the kidneys.
[68] 2018 CHOP 50:CAKUT TL+SVM $ACC:0.87\pm 2.1$ Improve CAKUT and kidney Propose a TL-based method and SVM model to extract Kidney
50:Healthy $AUC:0.92\pm 0.7$ diagnosis in children features from US kidney images to improve the CAKUT diagnosis.
Abbreviation: Kind of segmented organ (KSO)
I-C Paper Organization
The structure of this document is outlined below. Section 2 examines the current landscape of TL-driven KC through a detailed classification scheme. Following this, Section 3 outlines the significant advancements in various applications of TL-enhanced KC, such as in the detection of cancer, analysis of CT scans, examination of genomic data, and the integration of multi-modal data. Discussions on the datasets used for KC and the primary benefits and drawbacks of employing TL in the context of KC are presented in Sections 4 and 5, respectively. Subsequently, Section 6 delves into the obstacles faced in KC applications utilizing TL and outlines potential areas for future investigation in Section 7. Conclusions are drawn in the final section, Section 8.
II Taxonomy of TL
TL is a method in machine learning (ML) that involves applying knowledge obtained from solving one problem to a different, yet related, problem. This part presents the mathematical framework underlying TL, beginning with the concepts of a ”domain” and a ”task.”
A domain $\mathcal{D}$ consists of two primary elements:
  • A feature space $\mathcal{X}$, encompassing all potential inputs.
  • A marginal probability distribution $P(X)$, with $X\in\mathcal{X}$, which describes the distribution of samples across the feature space.
A task $\mathcal{T}$ is distinguished by:
  • A label space $\mathcal{Y}$.
  • A goal-oriented predictive function $f(\cdot)$ that associates an input $x\in\mathcal{X}$ with an output $y\in\mathcal{Y}$. This function $f$ is generally unknown and represents the target of learning algorithms to estimate.
In TL, we deal with a source domain $\mathcal{D}_{S}$ and a target domain $\mathcal{D}_{T}$, along with their corresponding tasks $\mathcal{T}_{S}$ and $\mathcal{T}_{T}$. The objective is to improve the learning of the target predictive function $f_{T}(\cdot)$ in $\mathcal{D}_{T}$ using the knowledge from $\mathcal{D}_{S}$ and $\mathcal{T}_{S}$.
In the context of TL, we engage with both a source domain $\mathcal{D}_{S}$ and a target domain $\mathcal{D}_{T}$, as well as their respective tasks $\mathcal{T}_{S}$ and $\mathcal{T}_{T}$. The aim is to enhance the learning of the target predictive function $f_{T}(\cdot)$ within $\mathcal{D}_{T}$ by leveraging the knowledge from $\mathcal{D}_{S}$ and $\mathcal{T}_{S}$.
The purpose of a TL approach is to utilize the source domain and task to improve the accuracy of the predictive function in the target domain. This involves adapting models, identifying shared feature representations, or employing other methods to make use of the similarities between the domains and tasks. Table III, summarizes the types of TL. Table IV provides a summary of various cutting-edge techniques that utilize TL types for KC, including the datasets and metrics employed.
TABLE III: Types of TL.
TL settings Source Target Tasks
Traditional ML/DL $D_{s}=D_{T}$ $D_{s}=D_{T}$ All
Inductive TL $D_{s}\cong D_{T}$ $D_{s}\neq D_{T}$ Classification and regression
Transductive TL $D_{s}\cong D_{T}$ $D_{s}=D_{T}$ Classification and regression
Unsupervised TL $D_{s}\neq D_{T}$ $D_{s}\neq D_{T}$ Clusterings and reduction
II-A Inductive TL
Inductive TL focuses on using the knowledge gained from the source domain and task to improve learning in a related but different target task. The key elements are defined as follows:
  • Let $\mathcal{D}_{S}=\{X_{S},P(X_{S})\}$ and $\mathcal{D}_{T}=\{X_{T},P(X_{T})\}$ be the source and target domains, respectively, where $X$ represents the input space and $P(X)$ the marginal probability distribution.
  • The tasks $\mathcal{T}_{S}=\{Y_{S},f_{S}(\cdot)\}$ and $\mathcal{T}_{T}=\{Y_{T},f_{T}(\cdot)\}$, with $Y$ being the output space and $f(\cdot)$ the target predictive function.
  • The goal is to improve $f_{T}$ by utilizing $f_{S}$, often requiring some labeled data in $\mathcal{D}_{T}$.
II-A1 Fine-tuning
During fine-tuning, the model’s weights, initially learned from the source domain, are adjusted to better suit the target task (Fig.4(a)). This adjustment is done by continuing the training process on the target dataset, often with a smaller learning rate to make subtle changes to the weights. Fine-tuning in this context allows the model to adapt the knowledge acquired from the source domain to the specifics of the target domain and task, making it a quintessential example of inductive TL.
In [69], the authors’ research highlights the significance of fine-tuning in TL for enhancing a pre-trained VGG19 and naïve inception model’s ability to detect kidney diseases from CT images. By adapting the DL model more specifically to this task, fine-tuning significantly boosts performance, achieving an impressive accuracy in identifying various kidney diseases.
The authors in [70], focused on exploring a variety of techniques to improve model performance on enhancing the efficiency of human kidney segmentation (HKS) using U-Net architectures for computer-aided detection (CAD) and diagnosis. Among these techniques, they particularly emphasized the importance of hyperparameter fine-tuning, alongside test time data augmentations, train time data augmentations, shuffling, randomization, and normalization. This advancement underscores the critical role of meticulous data preprocessing and model optimization techniques, such as hyperparameter fine-tuning, in refining the accuracy and efficiency of DL models for precise medical image segmentation tasks.
In the study referenced by [71], the researchers introduce a smart model for classification and prediction. This model employs an enhanced deep belief network for the task of predicting diseases related to the kidneys. It incorporates fine-tuning through the back-propagation method to adjust all network parameters optimally using the up-down algorithm. This algorithm is a contrastive version of the wake-sleep algorithm, which is applied for training deep belief network with labeled datasets. The assessment of this model demonstrates its effectiveness in the early detection of chronic kidney disease and its stages, thus contributing to slowing down kidney damage progression.
In the research cited by [72], the authors introduce a DL-driven segmentation framework that emphasizes the critical role of fine-tuning. The framework operates through a two-phase process, starting with broad segmentation and moving towards more precise segmentation, leveraging advanced DL techniques. A key innovation of this framework is the adoption of a novel loss function designed to incorporate the mean and variance of pixel values specific to kidneys and renal tumors, thereby allowing for meticulous refinement. The challenge of limited data is addressed by pre-training the model on images from abdominal datasets. This initial step equips the model with the capability to discern fundamental characteristics from abdominal images, facilitated by TL. This approach underscores the importance of fine-tuning in enhancing the model’s accuracy and applicability in medical imaging.
In [53], the fine-tuning is used to involves adapting a DL model, specifically a 3D CNN called NephNet3D, to improve its ability to classify cell types in human kidney tissue using 3D nuclear staining. The fine-tuning significantly improved the model’s balanced accuracy on the new specimen’s data, demonstrating that fine-tuning NephNet3D with a small subset of labeled data from a new specimen significantly enhances its predictive performance on that specimen.
Fig.4(a) illustrates two methods of fine-tuning CNN: (i) to fine-tune the weights of the whole network and (ii) to fine-tune some certain layers. Fig.4(b) showcases a CNN instance of deep network adaptation, which modifies the distribution in the full connection layer by measuring the domain distance.
Figure 4: Example of TL models used in KC: (a) Fine-tuning, and (b) DA.
II-A2 Multi-Task Learning (MTL)
Multi-task Learning aims at simultaneously learning multiple related tasks to improve generalization. It is mathematically described as:
  • Given a set of tasks $\{\mathcal{T}_{1},\mathcal{T}_{2},\ldots,\mathcal{T}_{N}\}$, the objective is to learn these tasks jointly by exploiting commonalities and differences.
  • This involves optimizing a shared representation $\Phi$ and task-specific functions
    $\{f_{T_{1}},f_{T_{2}},\ldots,f_{T_{N}}\}$.
  • The objective function is
    $\min_{\Phi,f_{T_{1}},\ldots,f_{T_{N}}}\sum_{i=1}^{N}\alpha_{i}\mathcal{L}_{T_{i}}(f_{T_{i}}(\Phi(X_{T_{i}})),Y_{T_{i}})$, where $\mathcal{L}_{T_{i}}$ is the loss function for task $i$.
The authors’ objective in [73] is to introduce an innovative semi-supervised multi-task learning method designed for forecasting the short-term progression of kidney disease using data from general practitioners’ electronic health records. They have shown that their methodology can: (i) track the progression of estimated glomerular filtration rate over time by enforcing a connection between sequential time periods, and (ii) harness valuable insights from patient data without labels, particularly when the data with labels are in limited supply.
In [74], the authors introduced a new multi-task learning approach aimed at assessing image quality, performing de-blurring, and classifying kidney diseases. The study analyzed immunofluorescence images from 1,608 patients, dividing them into 1,289 for training and 319 for testing. This innovative technique was not only effective in diagnosing four different kidney diseases using blurred immunofluorescence images, but it also demonstrated strong capabilities in two additional tasks: evaluating image quality and removing blur from images.
In [75], the authors introduce a multi-task 3D CNN designed for segmenting autosomal dominant polycystic kidney disease, which is identified by the continuous enlargement of renal cysts and stands as the most common severe hereditary kidney disorder. They successfully attained a high average dice score through this approach.
II-B Transductive TL
The objective of transductive TL is to predict the labels $\{y_{j}^{T}\}_{j=1}^{n_{T}}$ for the instances in the target domain $\mathcal{D}_{T}$, leveraging the knowledge from the source domain $\mathcal{D}_{S}$.
  • A source domain $\mathcal{D}_{S}=\{(\mathbf{x}_{i}^{S},y_{i}^{S})\}_{i=1}^{n_{S}}$ with $n_{S}$ labeled instances, where $\mathbf{x}_{i}^{S}\in\mathbb{R}^{d}$ represents the feature vector of the $i$-th instance and $y_{i}^{S}\in\mathcal{Y}$ represents its label.
  • A target domain $\mathcal{D}_{T}=\{\mathbf{x}_{j}^{T}\}_{j=1}^{n_{T}}$ with $n_{T}$ unlabeled instances, where $\mathbf{x}_{j}^{T}\in\mathbb{R}^{d}$.
Transductive TL can be further categorized based on the strategies employed to bridge the gap between the source and target domains:
II-B1 Domain Adaptation
It focuses on adapting the source domain model to work effectively on the target domain by minimizing the domain shift. Techniques include feature representation alignment, domain-invariant feature learning, and discrepancy minimization between source and target domain distributions. The aim of DA is to minimize the domain shift between the source domain $\mathcal{D}_{S}$ and the target domain $\mathcal{D}_{T}$ (Fig.4(b)) [76].
  • Feature representation alignment: Aim to find a transformation $\phi$ that aligns the feature distributions of the source and target domains by minimizing a distance metric $d(\phi(\mathcal{D}_{S}),\phi(\mathcal{D}_{T}))$, such as the maximum mean discrepancy [77].
  • Domain-invariant feature learning: Learn a feature representation $\phi$ that is invariant to the change in domains, optimizing $\phi$ such that the predictive model $f(\phi(\cdot))$ achieves low error on both $\mathcal{D}_{S}$ and $\mathcal{D}_{T}$ [78].
  • Discrepancy minimization: Directly minimize the discrepancy between the source and target domain distributions using techniques like adversarial training, where a discriminator is trained to distinguish between the source and target domain features while the feature extractor $\phi$ is trained to fool the discriminator [79].
Instance weighting, feature transformation, and kernel methods are key techniques used in the DA process, aimed at addressing the issue of domain shift—where the data distribution in the source domain differs from that in the target domain [80]. These techniques facilitate the transfer of knowledge from a source domain to a target domain, helping to improve the performance of ML models on the target domain despite differences in data distribution [81].
  • Instance weighting: It assigns weights to instances in the source domain based on their relevance to the target domain. The goal is to give higher importance to source instances that are more similar to the target domain, thereby reducing the bias towards the source domain distribution [82]. Typically, given a source instance $(x_{i}^{S},y_{i}^{S})$, assign a weight $w_{i}$ based on its similarity to the target domain. The weighted loss for the source domain is then $\sum_{i=1}^{n_{S}}w_{i}\ell(f(x_{i}^{S}),y_{i}^{S})$, where $\ell$ is a loss function [83].
  • Feature transformation: It transforms the feature space to find a common subspace where the source and target domain data can be directly compared or combined. Methods include subspace alignment, manifold learning, and DL approaches that learn domain-agnostic feature representations [84].
    • Subspace alignment: Find transformations $\phi_{S}$ and $\phi_{T}$ for the source and target domains, respectively, that align their subspaces, minimizing $\|\phi_{S}(\mathcal{D}_{S})-\phi_{T}(\mathcal{D}_{T})\|$ [85].
    • Manifold learning: Discover a low-dimensional manifold that both domains can be projected onto, learning a mapping $\phi$ that captures the intrinsic geometry of the data [86].
    • DL approaches: Use neural networks to learn a domain-agnostic feature representation, possibly with domain-adversarial components [87].
  • Kernel methods: They employ kernel functions to map the original feature space into a higher-dimensional space, where the source and target domains are more comparable. This approach facilitates finding a common representation that reduces the distributional differences between the domains [88]. Given a kernel function $k(\cdot,\cdot)$, transform instances $x_{i}^{S}$ and $x_{j}^{T}$ into a higher-dimensional space via $\phi$, where $\phi$ is implicitly defined by $k$. The goal is to learn a predictive model in this space that reduces the domain discrepancy [89].
These strategies are not mutually exclusive and can be combined to develop more sophisticated transductive TL models that effectively utilize the knowledge from the source domain to make accurate predictions in the target domain.
In [90], the authors propose a two-stage source-free unsupervised domain-adaptation framework for cross-modality abdominal multi-organ segmentations. The proposed method only requires a well-trained source model and unlabeled target dataset and does not require access to the source data. The experimental results highlight that this approach achieves satisfactory performance for adaptation of a labeled CT dataset to an unlabeled MRI dataset on abdominal organ segmentations including the right kidney and left kidney.
In [91], the authors present a new DA approach called dimension-based disentangled dilated DA to disentangle the storage locations between the features to tackle the problem of domain shift for medical image segmentation tasks without the annotations of the target domain. They use adaptive instance normalization to encourage the content information to be stored in the spatial dimension, and the style information to be stored in the channel dimension. They validate the proposed method for cross-modality medical image segmentation tasks on public datasets such as liver, right kidney, left kidney and spleen.
II-B2 Domain Generalization
Domain generalization focuses on the challenge of developing models that can generalize well to unseen domains. In domain generalization, a model is trained on data from multiple source domains to perform effectively on data from target domains that were not encountered during training. The core objective is to learn a domain-invariant representation that captures the underlying structure of the data, enabling the model to make accurate predictions across different, unseen domains [92]. In domain generalization, the goal is to learn a model $f$ that generalizes well to any target domain $D_{T}$, given multiple source domains $\{D_{1},D_{2},\ldots,D_{K}\}$ [93]:
$\min_{f}\sum_{k=1}^{K}\mathcal{L}(f(X_{k}),Y_{k})+\lambda\Omega(f)$ (1)
In [94], the authors present an innovative approach utilizing neural style transfer to reduce model bias towards texture and intensity, aiming to prioritize shape characteristics. They conduct experiments using 200 T2-weighted MR images, half with fat-saturation and half without, to segment kidneys in patients with polycystic kidney disease. The results demonstrate a significant improvement in segmentation performance, indicating better alignment with actual kidney shapes.
In [95], the authors address the challenge of single-source domain generalization in medical image segmentation, specifically noting the issue of domain shifts common in clinical image datasets. Traditional approaches, relying on global or random data augmentation, often produce samples that lack diversity and informativeness, inadequately representing the potential target domain distribution. To tackle this, the authors propose a novel data augmentation strategy for domain generalization, inspired by the inherent class-level representation invariance and style mutability of medical images. They suggest that unseen target data could be represented as a linear combination of class-specific random variables, each adhering to a location-scale distribution. This concept is operationalized through the use of constrained bézier transformation for augmenting data at both global and class-specific levels, significantly enhancing augmentation diversity. Additionally, a saliency-balancing fusion mechanism is introduced to increase the informativeness of augmented samples by utilizing gradient information for orientation and magnitude guidance. The authors theoretically demonstrate that their augmentation strategy effectively limits the generalization risk for unseen target domains. Their saliency-balancing location-scale augmentation method notably outperforms existing state-of-the-art techniques in two single-source domain generalization tasks, underscoring the potential of their approach to improve kidney and other medical image segmentation tasks by addressing the challenges posed by domain shifts.
II-B3 Self-Taught Learning
In self-taught learning, the model learns from unlabeled data that is not necessarily from the same distribution as the labeled data it was initially trained on. The key idea behind self-taught learning is to leverage large amounts of unlabeled data to improve the learning algorithm’s performance on a related task for which labeled data is scarce [96]. The process typically involves two main steps: first, extracting useful features or representations from the unlabeled data, and then, using these features to enhance the learning process on the labeled dataset. Self-taught learning enables the model to generalize better to new, unseen data by exploiting the underlying structure shared between the labeled and unlabeled datasets, even when they come from different domains. This approach is particularly valuable in scenarios where acquiring labeled data is expensive or labor-intensive, allowing for more efficient use of available data resources [97].
$\min_{f}\mathcal{L}(f(X_{S}),Y_{S})+\alpha\mathcal{L}_{AE}(X_{T})$ (2)
II-B4 Few-Shot Learning
Few-shot learning is a subset of transductive TL that focuses on the ability of a model to learn and make accurate predictions from a very limited number of training examples. The primary challenge in few-shot learning is to design algorithms that can quickly adapt to new tasks or classes using only a few labeled samples, often referred to as ”shots.” This approach is crucial in scenarios where collecting large amounts of labeled data is impractical or impossible, such as in medical diagnosis, where rare diseases may only have a handful of known cases. Few-shot learning leverages prior knowledge, gained from related tasks or a broader dataset, to infer the new task’s structure with minimal data. Typically, In few-shot learning, the model learns from a few examples in the target domain through a meta-learning approach:
$\min_{f}\sum_{\tau\in T}\mathcal{L}(f_{\theta+\Delta\theta_{\tau}}(X_{\tau}),Y_{\tau})$ (3)
This involves techniques such as meta-learning, where the model learns learning strategies from multiple tasks, or embedding learning, where data is transformed into a space where tasks with little data can be more easily learned. Few-shot learning aims to mimic human learning’s efficiency and flexibility, enabling models to generalize well from limited information.
In [98], the authors address the challenge of domain adaptation in DL for medical imaging, specifically in classifying kidney stones from endoscopic images taken under various conditions. They introduce a novel approach combining few-shot learning with meta-learning. Initially, a self-supervised learning phase pre-trains the model to improve feature generalization across domains. Subsequently, meta-learning fine-tunes these features specifically for the domain of kidney stones, adapting the model to new, unseen domains effectively. This method significantly enhances the model’s generalization ability and demonstrates a promising direction for overcoming domain adaptation challenges in medical imaging applications.
The paper [99] introduces the affinity network model, an innovative approach to few-shot learning, which excels in learning from a minimal set of examples and generalizing effectively. Its core is formed by layers of k-nearest-neighbor attention pooling, an advancement of the graph attention model, versatile enough to handle various object sets, with or without an underlying graph structure. This k-nearest-neighbor attention pooling is designed to be integrated seamlessly into any neural network, akin to how convolutional layers are used. A particular subset, the feature attention layer, is adept at highlighting critical features for classification. Through testing on both artificial and real-world genomic data from the the cancer genome atlas (TCGA) database, the affinity network has demonstrated superior generalization abilities compared to conventional CNN models, especially when trained on limited KC datasets.
In this study [100], the authors present a 3D few-shot segmentation framework tailored for precise segmentation of conditions such as KC, despite the scarcity of target organ annotation in training samples. A network resembling U-Net is developed to infer segmentation by understanding the relationship between 2D slices from support data and a query image, incorporating a bidirectional gated recurrent unit to ensure consistency in the encoded features among adjacent slices. Furthermore, they propose a TL approach to adjust the model to the specificities of the target image and organ by pre-testing updates using arbitrarily selected support and query samples. This model significantly outperforms existing few-shot segmentation models and rivals fully supervised models trained on more extensive target data.
In [101], the authors present an innovative approach for few-shot segmentation in medical imaging, specifically targeting left and right KC, to facilitate rapid adaptation of segmentation models to novel, previously unseen classes with minimal training data. They developed a segmentation framework based on episodic training within a CNN structure. Leveraging the natural spatial consistency found in medical images, a global correlation module was devised to determine the correlation between support and query images, integrating this module directly into the network. This technique is designed to enhance the network’s discriminative power, clustering features of the same class closer together while distinctly separating those of different organs. The efficacy of this method was validated using anatomical images from both CT and MRI scans, including cases related to KC.
II-C Unsupervised TL
Unsupervised TL focuses on transferring knowledge from a source domain to a target domain, where both domains lack labeled data, or the target domain lacks labeled data but the source domain may have some. The mathematical background of unsupervised TL (UTL) involves techniques and models that aim to learn useful representations or patterns that can be applied to new, unseen data. While there isn’t a single mathematical framework that encompasses all of UTL, several key concepts are commonly employed, including DA, representation learning, clustering and density estimation and autoencoders and generative models
II-C1 Representation Learning
The goal of representation learning is to find a transformation of the data that makes it easier to extract useful information for classification, regression, or other tasks. Mathematically, this involves learning a function $f:\mathcal{X}\rightarrow\mathcal{Z}$, where $\mathcal{X}$ is the input space and $\mathcal{Z}$ is a feature space that ideally captures the underlying factors of variation in the data.
In [102], the authors detail a self-supervised learning technique aimed at delineating the morphological features of molecular structures in microscopy images, critical for diagnosing disease phenotypes. They adeptly encode images of the slit diaphragm into a high-dimensional space without reliance on conventional segmentation or quantification methods. This process reveals a correlation between the self-supervised learning-generated embeddings and the slit diaphragm’s morphometric parameters, validated by established methodologies. Additionally, these embeddings mirror albuminuria levels, a key marker of kidney disease progression, and can distinguish between healthy and diseased mouse models with precision akin to traditional analyses. This self-supervised learning method facilitates insightful, label-free analysis of biomedical images, setting the stage for further refined, task-specific studies.
In [103], the authors tap into electronic health records to investigate the various patient subgroups within chronic kidney disease, applying the ConvAE open-source framework to conduct unsupervised DL on a dataset comprising 2.2 million U.S. patients from the OPTUM® EHR database. Through this methodology, they employ representation learning to create numerical representations of patients, enabling the identification of chronic kidney disease subtypes, associated comorbidities, and demographics of rare disease groups. Additionally, by modifying an existing algorithm for assessing disease severity, they pinpoint patient subgroups prone to rapid chronic kidney disease progression. This innovative combination of representation learning and data-driven analysis offers a novel perspective on chronic kidney disease diversity and identifies patients at an increased risk of progressing to end-stage kidney disease, underscoring the importance of early intervention.
II-C2 Clustering and Density Estimation
Unsupervised learning often involves clustering, where the goal is to group the data into clusters that represent underlying patterns or structures. Mathematically, this can be formulated as optimizing an objective function that measures the quality of the clustering. For example, the k-means clustering algorithm aims to minimize the within-cluster sum of squares:
$\min_{\{\mathbf{c}_{k}\}}\sum_{k=1}^{K}\sum_{\mathbf{x}_{i}\in C_{k}}\|\mathbf{x}_{i}-\mathbf{c}_{k}\|^{2}$
where $C_{k}$ is the set of points assigned to cluster $k$ and $\mathbf{c}_{k}$ is the centroid of cluster $k$.
Density estimation, such as using gaussian mixture models, involves estimating the probability distribution that generated the observed data. This can be particularly useful for capturing the complex distributions of data in both source and target domains.
In [104], the authors detail a method to assess kidney function loss after renal transplants by monitoring monthly estimated glomerular filtration rate changes over two years. The approach involves principal component analysis to categorize the estimated glomerular filtration rate data into two clusters, followed by the use of kernel density estimation to analyze the distribution of estimated glomerular filtration rate values within these clusters. This reveals distinct patterns in kidney function post-transplant, important for patient monitoring and treatment adjustments.
In [105], digital image processing techniques were applied to 3D CT scans to identify and diagnose various kidney disorders, including stones, cancer (tumors), cysts, and renal fibrosis, alongside healthy kidney tissue. Utilizing k-means clustering and co-occurrence matrices for segmentation and texture analysis, this method effectively distinguishes between healthy and diseased tissues, enabling the detection of tumors at very early stages. Additionally, density slicing transforms grayscale images into color-coded images based on texture density, aiding in the visualization of different tissue conditions. The analysis revealed that KC tissues exhibit distinct statistical properties like higher standard deviation and mean values, contrasting significantly from other conditions, thus facilitating early diagnosis through detailed tissue characterization.
This study [106], addresses chronic kidney disease, characterized by a progressive decline in kidney function, commonly caused by diabetes and high blood pressure. It aims to enhance diagnosis accuracy and efficiency by applying clustering algorithms to patient test data. The research focuses on categorizing chronic kidney disease stages by severity using algorithms such as k-means, k-medoids, and Fuzzy C Means. Findings indicate that the Fuzzy C Means algorithm outperforms others in distinguishing between different chronic kidney disease stages, thereby promising quicker and more precise diagnoses.
In this study [107], examines the link between chronic kidney disease and cardiovascular risk factors within nigerian populations, specifically addressing the notable lack of research on the clustering of these conditions in black homogeneous groups. Conducted in south-west nigeria, this research utilized a modified world health organization questionnaire for gathering a range of data and performed biochemical analysis on plasma samples from 1084 participants. Findings revealed a 14.2% prevalence of stage 3 chronic kidney disease, with an increase in hypertension and low HDL-C levels associated with greater clustering of chronic kidney disease and cardiovascular risk factors. The data indicated an inverse relationship between the estimated glomerular filtration rate and the aggregation of risk factors, highlighting the significant linkage between chronic kidney disease, proteinuria, and the clustering of cardiovascular risk factors. This study emphasizes the critical need for chronic kidney disease screening in individuals showing clusters of cardiovascular risk factors and suggests the importance of integrated management for these health issues.
II-C3 Autoencoders and Generative Models
Autoencoders learn a compressed, encoded representation of data, aiming to capture its most salient features. The autoencoder consists of an encoder function $f(\mathbf{x})$ that maps an input to a hidden representation, and a decoder function $g(\mathbf{h})$ that reconstructs the input from the hidden representation. The learning objective is to minimize the reconstruction error:
$\min_{f,g}\sum_{\mathbf{x}\in\mathcal{X}}\|\mathbf{x}-g(f(\mathbf{x}))\|^{2}$
Generative models, like variational autoencoders and generative adversarial networks, aim to model the distribution of data in such a way that new, synthetic data points can be generated. These models are powerful tools for unsupervised learning and TL, as they can capture complex data distributions.
The study [108] introduces a DL framework designed to diagnose cancer by leveraging pre-trained autoencoder technologies for effective feature extraction from complex, high-dimensional gene data, with a focus on KC. By fine-tuning the model for specific tasks, the framework uses neural network classifiers to predict clinical information accurately. Notably, the model significantly outperformed traditional techniques, achieving high micro and macro f1-score (F1) across gender, race, and sample type predictions. Specifically, the introduction of a conditional variational mutation autoencoder notably enhanced race prediction accuracy. The findings underscore the framework’s potential in improving prognosis, prevention, and early detection of KC.
TABLE IV: A summary of TL types for KC.
Ref/Year Model-based TL TL type KC task Datasets Image features Metrics
[69] (2022) VGG19 and Naïve inception Inductive Fine-tuning Detection Private data 4000, CT ACC
[70] (2022) U-Net Inductive Fine-tuning Segmentation HuBMAP 15, CT Dice score
[72] (2024) U-Net Inductive Fine-tuning Segmentation Private data CT -
[53] (2023) 3D-MS-RFCNN Inductive Fine-tuning Segmentation KiTS 1059, 512×512, CT ACC and dice score
[73] (2021) - Inductive MTL Prediction Private data CT Recall
[74] (2021) CNN Inductive MTL Classification Private data 1608, CT ACC and AUC
[75] (2018) 3D CNN Inductive MTL Segmentation Private data 203, CT Dice score
[90] (2022) - Transductive DA Segmentation Private data 30, CT and 20, MRI Dice score
[91] (2023) - Transductive DA Segmentation Private data 20, CT and 20, MRI Average
[94] (2021) CNN Transductive DG Segmentation Private data 200, MRI Dice score
[95] (2023) - Transductive DG Segmentation Private data MRI Dice score
[98] (2022) DL Transductive FSL Classification Private data 600, CT ACC
[99] (2019) KNN Transductive FSL Prediction GDCD 654 samples ACC
[100] (2021) U-Net Transductive FSL Segmentation BCV 30, 3D, CT Dice score
[101] (2022) CNN Transductive FSL Segmentation Private data 3D, 20, MRI, 30, CT Dice score
[102] (2024) CNN Unsupervised RL Segmentation Private data 288, CT ACC, F1
[103] (2022) CNN Unsupervised RL Classification OPTUM®EHR US ACC
[105] (2023) K-Mean Unsupervised CDE Classification - - Mean
[106] (2017) K-Mean Unsupervised CDE Prediction UCIMLR 400 instances ACC, SEN, SPE
[108] (2022) DL Unsupervised Autoencoders Classification - - F1
III Applications of TL in KC
TL has been applied in the field of KC diagnosis and prognosis using medical imaging and genomic data. Some applications include:
III-A Cancer Detection
TL is a method in DL that utilizes knowledge acquired from one task to enhance performance on a novel, but related, task, as outlined in [109]. Within the scope of analyzing medical images, TL enables the training of models on extensive datasets of analogous images, like radiology scans, before refining these models on a smaller set of medical images designated for a specific purpose, such as the detection of KC. To deploy TL for KC detection, the initial step involves gathering a comprehensive dataset of medical scans, such as CT or MRI images, which includes both healthy kidneys and those with cancerous growths. This collection serves as the foundation for pre-training a CNN to differentiate between normal and cancerous images. After the CNN has been pre-trained on the extensive dataset, it is then fine-tuned using a more focused collection of kidney images, precisely for KC detection. This fine-tuning adjusts the CNN’s pre-trained weights to refine its performance for this specific detection task. Throughout the fine-tuning phase, the CNN hones in on distinct characteristics within the kidney images that signal the presence of cancerous tumors. This process benefits from the insights gained from the larger dataset of similar images, allowing the CNN to learn how to identify KC in the smaller dataset with increased speed and accuracy [110].
III-B CT Scan Analysis
TL is a prominent strategy in DL, facilitating the adaptation of pre-trained models to novel tasks. This method has been particularly effective in the medical imaging domain for tasks such as the automatic segmentation and classification of kidney tumors from CT scans. TL operates on the principle of leveraging the insights acquired from a model trained on a related issue trained on a similar dataset. This strategy is advantageous because it conserves time and resources that would otherwise be expended on initiating training from the ground up, while also enhancing accuracy through the transference of knowledge from the pre-trained model. Specifically, in the analysis of CT scans, TL has proven beneficial in refining pre-trained DL models for more precise and efficient identification and categorization of kidney tumors, as reported in [111, 112].
III-C Genomic Data Analysis
TL stands as a formidable method within ML, permitting the application of insights acquired from addressing one issue to enhance outcomes on a closely related issue. This methodology has seen use in genomics for forecasting the prognosis of KC patients through gene expression data and distinguishing KC’s molecular subtypes. The core concept of TL is to harness the knowledge from a pre-existing solution to a comparable challenge, specifically, utilizing a model pre-trained on extensive datasets, to bolster the new model’s efficacy on a pertinent task. The adoption of TL allows for a conservation of time and resources that would be otherwise spent on building a model from the ground up, simultaneously achieving superior results by drawing upon the learnings from prior endeavors. Within the realm of genomic data analysis concerning KC, TL has facilitated more accurate and efficient prognostic predictions and molecular subtype identification, significantly refining the diagnosis and treatment strategies for this ailment [113, 114, 115].
III-D Multi-Modal Data Fusion
TL has been instrumental in enhancing the precision of KC diagnosis and prognosis by consolidating diverse data sources, including CT scans, MRI images, and gene expression data. This approach allows for a comprehensive analysis by leveraging the strengths of each data type. For instance, imaging data like CT and MRI scans offer visual insights into the physical state of the kidneys, identifying tumors’ size, shape, and location. On the other hand, gene expression data provides a molecular perspective, offering clues about the tumor’s aggressiveness and potential response to treatment. By applying TL, models pre-trained on large datasets from one domain (e.g., radiology imaging) can be fine-tuned with data from another domain (e.g., genomic data), facilitating a multi-faceted understanding of KC. This integration enables the models to learn from the complex interplay between the imaging characteristics of tumors and their molecular profiles, leading to more accurate and personalized diagnostic and prognostic assessments. The ability to draw upon pre-existing knowledge from various domains not only saves significant time and resources but also enhances the models’ predictive capabilities, making TL a powerful tool in the fight against KC [116, 117]. Fig.5 illustrate the applications of TL in KC.
Figure 5: The applications of TL for KC
IV Existing Datasets
The efficacy of CAD systems is significantly dependent on the dataset they are trained on, whether these systems employ TL techniques or other methodologies. The volume and integrity of the dataset not only affect its quality but also the performance of the resultant models. Yet, the acquisition of extensive medical imaging data presents challenges due to the intricate nature of imaging techniques and potential risks to patients. Moreover, the process of annotating such data demands the expertise of skilled radiologists to ensure precision, making high-quality, large-scale annotated datasets rare. In response to this limitation, research into novel data augmentation strategies, including photometric and geometric alterations, is underway to enrich existing datasets, thereby alleviating the issue of data scarcity. Enhancing the dataset’s quality via advanced imaging technique is another strategy to bolster the performance of models based on TL. These initiatives pave the way for enhanced model outcomes through the use of enriched data. There are several key datasets accessible to the public that offer insights into kidney tumors, namely: (i) The kidney tumor segmentation 19 (KiTS19) challenge dataset, offering CT scans for tumor segmentation algorithm assessment. (ii) The national cancer institute’s kidney renal clear cell carcinoma dataset, with gene expression and clinical details pertinent to clear cell renal cell carcinoma, the prevalent form of KC. (iii) The TCGA kidney renal papillary cell carcinoma dataset, which provides genomic and clinical data for patients with papillary renal cell carcinoma. (iv) The dataset from radboud university medical center focusing on kidney tumors, featuring CT scans alongside segmentation masks for evaluating kidney tumor segmentation techniques. Each dataset is distinct in terms of the data it encompasses, the patient count, and the imaging techniques employed, as summarized in Table. V, which centers on datasets related to kidney tumors.
TABLE V: Summary of datasets-based kidney tumor.
Ref Datasets RR Applications The statistics of dataset
[118] UCIMLR [119], 2018 Classify CKDD using SVM 400 cases; 25 attributes
[120], 2019 Diagnosis of CKDD using ANN and SVM 400 patients; 24 attributes
[121], 2020 Classify CKDD using MKSVM and FFOA 400 instances; 25 attributes
[122], 2021 Predict CKDD using TL 300 patients, 210 labels
[123], 2022 Kidney segmentation using R-CNN -
[124] TCGA [125], 2018 Distinction between RCC using ANN and SVM 26 patients
[126], 2019 Prediction the mutation in RCC using ANN and RF. 267 patients
[127], 2020 Differentiating high- from low-grade RCC using TL. 390 patients
[128], 2021 Integrate CT and genomics features for RCC using SVM-RFE 205 patients; 107 features
[129], 2022 Classify genomic characteristics in RCC using ML. 575 cases
[130] KiTS19 [63], 2019 Segmentation of kidneys from CT images of kidney stone patients. 153 patients
[26], 2020 Delimit the kidneys in CT images. 210 patients
[122], 2021 Kidney segmentation 300 patients; 210 labels
[131], 2022 Segment kidney and tumor regions. 20 patients, 210 images
[132] IRCAD [133], 2019 Renal segmentation. 363+128 patients
[134] MICCAI [135], 2019 Kideny segmentation. 210 patients
Abbreviation: Relevant references (RR),
V Advantages and Limitations of TL in KC
TL can have several advantages in the context of KC: Securing annotated data in the healthcare sector is a significant challenge, often requiring considerable effort. TL plays a pivotal role in this context by facilitating the use of existing pre-trained models. These models, which have been educated on large and universally accessible datasets, can be adapted to develop specialized models for diagnosing specific conditions like KC. The advantages of applying pre-trained models are manifold. Firstly, these models, having been exposed to vast amounts of data, have the capability to significantly enhance performance on smaller, task-specific datasets, particularly in diagnosing KC. This approach not only leads to considerable savings in time and financial resources but also accelerates the training process. This is because fine-tuning a pre-trained model for a new task is much faster than starting training from scratch. Furthermore, TL reduces the need for extensive data and computational power for training models on specific tasks. In the realm of medical imaging, where missing or scarce data can pose significant challenges, especially for rare diseases like KC, TL offers a valuable solution by employing models that have been pre-trained on extensive datasets. This method effectively bridges the data gap, providing a way to address the issues of scarce data and domain shift that are prevalent due to the high variability in medical imaging data, making it especially beneficial for scenarios involving small datasets.
In addition there are several limitations to using TL for KC diagnosis:
The application of TL in diagnosing KC presents several challenges and considerations. One primary challenge is domain adaptation, which necessitates identifying a pre-trained model trained on a domain or task closely related to KC diagnosis. The effectiveness of TL hinges on the alignment between the original and target domains and tasks, with discrepancies potentially undermining the model’s ability to transfer knowledge effectively. Moreover, the risk of overfitting is heightened when fine-tuning pre-trained models on limited datasets, leading to models that perform poorly on new or unseen data. Fine-tuning entails careful adjustments of model layers, learning rates, and regularization methods to avoid such overfitting. Data scarcity for the new task can further complicate fine-tuning, potentially resulting in suboptimal performance. Additionally, TL can be computationally demanding, especially with large pre-trained models, posing scalability challenges. Interpretability issues also arise, as understanding the decision-making process of a model trained on a different task can be complex. The biases in the datasets used for training pre-trained models may carry over, affecting performance on certain data subsets. Moreover, these models, being trained on specific datasets, might not generalize well across different datasets, especially when data distributions vary significantly. Privacy concerns are another critical consideration, particularly when sensitive medical data must be shared with third parties for model training. The architecture of the pre-trained model also plays a crucial role in the success of TL ; an ill-suited architecture can hinder effective fine-tuning. Lastly, the substantial computational resources required for TL can make the use of pre-trained models costly, especially when employing multiple models or conducting several rounds of fine-tuning. These challenges underscore the complexity and nuanced considerations involved in applying TL for medical diagnosis tasks such as KC. Table. VI, provides an overview of the benefits and drawbacks of TL-based kidney tumor techniques.
TABLE VI: Summary of advantages and limitations of TL-based kidney tumor techniques.
Work Year LM MPS Dim Metrics Advantages/Limitations
[136] 2022 CNN+TL Segmentation and classification 2D DC, ACC, Pr, SEN, F1, MSE Noise problem, the substantial size of WSIs and the computational costs
[137] 2022 CNN+TL Segmentation and classification 2D Pr, ACC, F1 The anticipated time was higher
[138] 2022 CNN+TL Segmentation and classification 2D ACC, SEN, SPE, SEN, F1, MSE Not accurate at higher levels and insufficient data
[139] 2022 SHUW+CNN Segmentation and classification 2D SEN, DC, SD Application to only one area and acceptable performance
[140] 2022 ResUNet and DenseUNet Segmentation and classification 2D DC Limited testing size, instability of data distribution
[141] 2022 CNN+TL Segmentation, and classification 3D DC, JI, FN, VS, ROC-AUC High complexity, the field of view not be sufficient
[142] 2022 Histo-Cloud+CNN Segmentation 2D Kappa, TPR, MCC The computational complexity
[143] 2022 TL+AE Segmentation, sample and feature space augmentation 2D ROC-AUC Latent representations, small data volume, imbalance datasets, improved predictive model
[144] 2021 U-Net Segmentation and classification 2D CrM, Pr, SEN, ACC, JI, DC Complex deep network and unnecessary computations have been reduced, low processing speed
[145] 2021 SVM Semantic segmentation, feature extraction and classification 2D ACC, SEN, Pr, F1 Motion artifacts, accuracy issue, limited datasets and
[146] 2020 TL+FCNN Segmentation 2D DC, F1 Reducing in the training time, segmentation accuracy and smaller datasets
Abbreviation: Learning model (LM), Main processing steps (MPS), Dimention (Dim),
VI Key Challenges
The studies reviewed in this article emphasize the effectiveness of TL for diagnosing KC, noting its computational efficiency and potential to outperform conventional ML and DL techniques. However, several key challenges still need to be addressed to enhance the performance of TL. This section aims to highlight the most significant unresolved issues and current concerns that are receiving considerable focus in the area of KC diagnosis:
VI-A Availability, Adapting and Adjusting Models
Data accessibility plays a pivotal role in crafting and applying TL models for diagnosing KC. Nevertheless, several challenges complicate data accessibility in this arena. A primary issue is the scarcity of annotated medical imaging data specific to KC. This data, essential for TL model training, comprises images marked with details indicating the presence or absence of diseases like KC, yet it’s often limited in both amount and quality. Another obstacle arises from the inconsistencies in imaging techniques and data collection methods among various hospitals and healthcare facilities, which can introduce irregularities in the data for TL models, adversely affecting their performance and precision. Moreover, concerns related to privacy and security pose significant restrictions on medical imaging data’s availability for TL, given the sensitive nature of patient information embedded in such data, complicating its distribution among distinct entities. Additionally, data bias presents a critical challenge; TL models trained on skewed data might exacerbate existing inequalities in medical diagnoses and treatments. For instance, a model trained predominantly on data from male subjects might underperform with female patient images, reflecting and potentially worsening disparities.
The primary obstacle in TL involves identifying an appropriately pre-trained model that has been trained on a domain or task closely related to the new task at hand. It’s critical to evaluate how similar the original (source) and the intended (target) domains are, as well as the resemblance between the tasks involved [147, 148].
Adjusting a pre-trained model for a specific task, known as fine-tuning, presents its own set of challenges. It necessitates a strategic selection of which layers to adjust, alongside the determination of an optimal learning rate and the application of appropriate regularization strategies to prevent the model from becoming too narrowly adapted to the task [149].
VI-B Data Scarcity
Data scarcity presents a significant challenge in the development and efficacy of kidney cancer diagnosis systems, impacting everything from image analysis to the training of deep learning models. Several studies have explored these challenges and proposed methods to mitigate the impact of limited data availability [150]. Peng et al. [151] explore how downsampling and compression affect renal image analysis. They found that data scarcity impacts diagnostic accuracy, highlighting the need for robust data handling techniques to ensure the integrity and usefulness of renal images in diagnosis. Similarly, Guo et al. [152] address the limited training data for kidney segmentation by using a cascaded convolutional neural network, demonstrating that innovative network architectures can partly overcome the challenges posed by sparse data.
In the context of clinical applications, Heller et al. [153] provide the kits19 dataset, which includes 300 kidney tumor cases with CT semantic segmentations and surgical outcomes, aiming to enhance research in kidney cancer by making rich datasets available. This effort underscores the critical role of accessible high-quality datasets in advancing early tumor detection and improving surgical outcomes. Flitcroft et al. [154] and Nassour et al. [155] both discuss the scarcity of reliable diagnostic biomarkers for kidney cancer, particularly in non-invasive strategies like urinary protein analysis and the evaluation of Lynch syndrome’s impact on kidney cancer. These studies highlight the gap in epidemiological data and the need for comprehensive datasets to develop effective diagnostic tools.
Technological advancements in handling data scarcity are also evident in deep learning applications. Alzu’bi et al. [156] and Bhattacharjee et al. [157] discuss the use of CNNs for kidney tumor classification and detection. They focus on optimizing deep learning approaches to perform effectively even with limited data, a crucial advancement given the rarity and variability of KC data. Lastly, Wang et al. [158] delve into the utilization of deep learning techniques for the imaging diagnosis of renal cell carcinoma. They emphasize how these advanced computational tools allow clinicians to diagnose and evaluate renal tumors more accurately and swiftly. One of the key strategies discussed is leveraging unlabeled data, which is a valuable solution for coping with data scarcity. By utilizing semi-supervised learning techniques, the potential of existing datasets can be maximized without the need for extensive labeled data.
VI-C Data Bias and Generalization
Data bias represents a significant obstacle in TL for KC diagnosis, impacting model accuracy and fairness. Bias can manifest in various forms, such as sampling bias, where training data may not accurately reflect the disease’s true distribution, potentially leading to population-specific biases [159]. Measurement bias skews data collection towards specific attributes, like age or gender, causing model predictions to favor these characteristics. Annotation bias arises from annotator prejudices, affecting the reliability of data labels [160]. To counteract these issues, TL models may utilize data augmentation to enrich training data variety and volume. Training models on data from diverse sources, like multiple hospitals, ensures a representative dataset of the disease’s distribution. Additionally, implementing fairness constraints and promoting algorithmic transparency can help mitigate bias, fostering the development of fair and unbiased models [161].
Generalization stands as a pivotal challenge in TL for diagnosing KC, denoting the model’s capacity to deliver accurate predictions on unseen data. Achieving robust generalization is essential, determining the model’s viability in practical scenarios [162]. A notable hindrance to strong generalization is the limited availability of annotated data. For many conditions, including KC, annotated data for training may be scarce, leading to a risk of overfitting, where the model overly specializes on training data, diminishing its performance on novel data. Additionally, data diversity issues, where training data may not accurately represent the disease’s distribution, can impair generalization [163]. For instance, data sourced from a singular hospital might not mirror the broader disease prevalence across different locales. To mitigate these issues, TL models for KC diagnosis may incorporate data augmentation strategies to broaden the training data’s size and variety. Training on data from varied sources, like multiple healthcare facilities, ensures a more representative dataset. Furthermore, employing regularization methods like dropout aids in curbing overfitting, enhancing the model’s generalization capabilities [164].
VI-D Privacy Concerns, Expertise Shortage and Accuracy Instability
Privacy concerns emerge prominently in developing and applying TL models for KC diagnosis due to the necessity of accessing sensitive medical information, triggering significant privacy and security considerations [159]. A key challenge is safeguarding medical data privacy and security, demanding stringent measures against unauthorized access and potential breaches. Managing medical data with utmost care to align with privacy regulations is crucial [37]. The annotation process also presents privacy challenges, necessitating judicious annotator selection and bias minimization efforts. Employing privacy-preserving methods, such as differential privacy, is essential to protect patient confidentiality during annotation. Additionally, securing TL models against unauthorized interventions and ensuring their resilience is paramount. Providing patients with transparent data usage information and granting them control over their data usage decisions is vital for maintaining trust and privacy integrity [165].
The creation and implementation of TL models for KC diagnosis necessitate profound expertise in medical imaging and ML, yet expertise scarcity poses several challenges [166]. The limited pool of experts hampers the resources and knowledge necessary for these models’ development and application. Interdisciplinary collaboration, essential between medical and ML experts, is often obstructed by differing methodologies and communication barriers [167]. Adhering to medical imaging and mML best practices and standards is imperative for the responsible development and deployment of TL models. Ensuring patient safety and efficacy of these models requires a deep understanding of patient needs and the development of models that positively impact patient outcomes.
Improving TL model performance often involves utilizing pre-trained models on extensive datasets, such as those trained on the ImageNet database. Nonetheless, limited data availability can lead to accuracy instability. Various regularization techniques, such as SPAR, have been explored to stabilize accuracy by moderating the target network’s outer layer weights with reference to a starting point [168]. An innovative regularization method, DL transfer with attention, was introduced, focusing on preserving the source model’s outer layer outputs rather than merely constraining the model’s weights. DL transfer with attention minimizes empirical loss and aligns the two models’ outer layer outputs by selectively constraining feature maps identified by a supervised learning-trained attention mechanism .
VI-E Knowledge Gain and Data Annotation
Evaluating the knowledge gained by the TL model and its performance outcomes is crucial. Consequently, considerable research has been directed towards establishing precise metrics for assessing the effectiveness of these models and the extent of knowledge they acquire, aiding in the elucidation of complex issues [161]. Proposed metrics include transmission error, transmission loss, transmission ratio, and in-range ratio. Nevertheless, the specificity of these metrics often falls short of expectations, highlighted by the ongoing uncertainty regarding outcomes when other TL-based models are employed, such as in KC detection scenarios, where model performance might be suboptimal.
Data annotation plays an essential role in developing TL models for diagnosing KC. Yet, this process faces several significant hurdles. A primary issue is the considerable cost and time required to annotate extensive datasets accurately. The process demands substantial expert input and resources, making it both time-intensive and costly. Another obstacle is the absence of standardized practices in data annotation, complicating the comparison and amalgamation of datasets from diverse origins and potentially diminishing the TL models’ impact. Ensuring the annotations’ quality and precision further necessitates meticulous validation and quality control, adding to the resource and time expenditure. Moreover, achieving representativeness in annotations, ensuring they accurately reflect the target demographic, involves careful annotator selection and concerted efforts to minimize bias throughout the annotation process.
VI-F Negative Transfer and Overfitting
Negative transfer in the context of learning about KC refers to the difficulties that individuals may experience when acquiring new knowledge or skills related to the condition and its treatment [169, 170]. (1) Overgeneralization: Individuals may apply learned information too broadly, which may result in an incorrect or incomplete understanding of the condition and its treatment. (2) Overreliance on prior knowledge: Individuals may rely too heavily on prior knowledge about cancer or other medical conditions, which may not be applicable to KC.(3) Incompatibility between old and new learning: Prior learning about cancer or other medical conditions may be incompatible with new information about KC, leading to confusion and hindering the acquisition of new knowledge. (4) Resistance to change: Individuals may be resistant to changing their existing beliefs and attitudes about cancer, leading to difficulty in adapting to new information about KC. (5) Misapplication of prior knowledge: Individuals may misapply prior knowledge, leading to an incorrect or inefficient understanding of the condition and its treatment.
Overfitting, especially in the domain of KC diagnosis, describes the situation where a predictive model becomes excessively tailored to the training dataset, undermining its capability to perform accurately on new, unseen data. This often results from training a model with an overly complex set of parameters, rendering it over-specialized and ineffective for general prediction or diagnosis purposes [171, 172]. In the realm of KC, overfitting may result in inaccurate diagnoses, an increase in false positives, or diminished disease detection sensitivity, posing significant risks to patient management and therapy. Implementing measures such as regularization to curb model complexity and employing cross-validation to evaluate model efficacy on novel data are critical strategies to prevent overfitting.
The study by [173] introduces innovative regularization techniques, including large margin loss, focal loss, adversarial training, mixup, and data augmentation. These methods aim to correct the logit shift in underrepresented classes and address class imbalance by scrutinizing network behavior, providing a robust approach to mitigate overfitting.
Furthermore, [174] details the use of Five-fold cross-validation as a means to assess the performance of AI algorithms, specifically in the segmentation process involving seed placement and bounding box techniques for lesion patch extraction, coupled with deep CNN for distinguishing clear CCRCC from renal oncocytoma.
VII New Perspectives and Current Trends
The results of previous research indicate that, despite having achieved high levels of accuracy, there is still room for improvement in the determination of KC. To that end, scientists propose using a less common set of algorithms that can aid in accurately identifying these cells, especially in their early stages. Enhancing the capability of models to accurately determine the presence of KC at various stages while preserving accuracy and improving unbiasedness is crucial. The combination of feature selection methods and DL models may prove helpful in this regard. A larger database is required, and the algorithm must be refined and optimized for other classification tasks related to KC detection. The data collected from various medical institutions and even available databases can be used to evaluate these strategies. There are several potential future directions for the use of TL in the context of KC diagnosis:
VII-A Multimodal Learning
Multimodal learning represents a strategy that integrates information from diverse sources or data types to enhance learning and decision-making processes (Fig.6). Within cancer diagnosis, this approach entails the amalgamation of data from varied origins, including medical imaging (such as MRI and CT scans), genomic data, and clinical records (like patient history and laboratory results). Merging multimodal learning with TL for cancer diagnosis involves training a comprehensive model on a dataset encompassing medical images, genomic information, and clinical data across multiple cancer varieties. Subsequently, TL techniques are employed to refine the model for diagnosing a particular cancer type. For instance, a DL model could initially be educated on a broad array of medical images from different cancers, with TL utilized to tailor the model for specific cancer types. Furthermore, the model might also undergo training with genomic and clinical data, integrating these disparate data types through multimodal learning methods to enhance the model’s precision and robustness. In essence, the application of multimodal learning in conjunction with TL for cancer diagnosis promises to elevate diagnostic accuracy, diminish instances of false positives and negatives, and facilitate earlier, more precise cancer detection.
Figure 6: Multimodal learning for KC
VII-B Federated Learning
Federated learning (FL) [175] is gaining momentum in the field of KC diagnosis, prediction, and segmentation, demonstrating substantial potential in enhancing accuracy and preserving data privacy across various healthcare environments. This collaborative approach enables the utilization of data from multiple institutions without compromising patient confidentiality, addressing both the variability and volume of medical data essential for effective machine-learning models. In the study by Vekaria et al. [176], the integration of FL with TL is explored as a novel approach to enhance the detection of kidney abnormalities from medical imaging. This research demonstrates how FL can facilitate collaborative learning across different healthcare institutions without necessitating direct data sharing. By doing so, it not only ensures data privacy but also leverages diverse datasets to improve diagnostic accuracy and model robustness. Gao et al. [177] highlight the effectiveness of a federated learning-driven collaborative diagnostic system for metastatic breast cancer. The ability of this system to improve generalizability and diagnostic accuracy without compromising patient privacy offers a compelling model for KC diagnostics, particularly in settings with diverse patient populations and imaging protocols. Similarly, the work by Subashchandrabose et al. [178] introduces an ensemble-based FL framework for the classification of lung cancer, which can be analogously applied to KC. This approach integrates multiple machine learning models trained on disparate datasets to enhance both accuracy and generalizability. The success of this method underscores the potential of FL in overcoming the limitations posed by single-institution datasets, suggesting a scalable model for KC diagnosis that respects patient privacy.
Further extending the application of FL, Ma et al. [179] develop an assisted diagnosis model for cancer recurrence using federated DL. This model not only predicts the recurrence of cancer but also identifies potential recurrence locations, offering a significant tool for post-operative care and monitoring in KC patients. By incorporating various physical examination indicators and leveraging a federated architecture, this approach enhances prediction accuracy while safeguarding sensitive patient data. Moving forward, Yaqoob et al. [180] focus on the implementation challenges of traditional machine learning in skin cancer but their insights are equally applicable to KC. They highlight the advantages of FL in building privacy-aware healthcare systems that comply with stringent data privacy regulations. This review provides a solid foundation for understanding how FL can be adapted for KC diagnosis and prediction, ensuring data privacy without compromising on diagnostic capabilities.
Tan et al. [181] and Li et al. [182] explore advanced FL frameworks that incorporate TL and data augmentation techniques like SMOTE to enhance the detection and classification capabilities in medical imaging. These studies illustrate the effectiveness of FL in handling imbalanced datasets, which is a common issue in KC diagnosis due to the variability of lesion appearances and stages.
A study by Kumbhare et al. [183] provides an insightful framework utilizing FL for breast cancer detection, which can be adapted for KC. This study outlines an innovative AI training method that allows hospitals to collectively harness the power of vast datasets from multiple sources to improve diagnosis accuracy. The approach not only respects patient privacy but also reduces processing times, indicating a promising avenue for KC applications where early detection and accurate segmentation are crucial. Almufareh et al. [184] demonstrate the use of FL in developing a highly accurate breast cancer detection model. By utilizing DNNs and data augmentation techniques within a federated framework, this study shows significant advancements in model performance. The methodology promises applicability to KC, where complex imaging features require sophisticated model tuning and privacy-preserving collaborative efforts. Jimenez et al. [185] explore the memory-aware curriculum learning in a federated setting to improve the classification of breast cancer across different clinical sites. This approach, focusing on optimizing the order and frequency of local updates, could be particularly beneficial for KC diagnosis, where data from multiple institutions vary significantly.
Lastly, Heidari et al. [186] describe the use of blockchain-based FL for detecting malignant lung nodules, showcasing the method’s robustness in preserving data integrity and anonymity across collaborating entities. This model’s success in maintaining privacy while achieving high accuracy underscores its potential adaptability to KC, where similar challenges in data variability and privacy concerns are prevalent.
VII-C Self-Paced Learning (SPL)
Self-paced learning (SPL) has emerged as a potent methodology for enhancing medical diagnostic systems, particularly in the challenging arena of kidney cancer diagnosis. This learning strategy prioritizes learning samples based on their complexity, adapting the training process to focus progressively on the more difficult or informative cases as the model’s capability improves. This adaptive learning approach can significantly benefit the diagnosis of kidney cancer, where data variability and imbalance often pose significant challenges. One significant application of self-paced learning in the medical field is addressed by Wang et al. (2020), who developed a self-paced learning sampling method specifically designed for imbalanced cancer classification. This method helps to reduce noise from imbalanced samples, improving the pre-diagnosis accuracy of cancer data, which is crucial for conditions like kidney cancer where early detection can significantly influence treatment outcomes [187]. Furthermore, the integration of SPL with federated learning, as discussed by Wang and Zhou, creates a privacy-preserving disease diagnosis framework [188]. This combination is particularly pertinent in decentralized healthcare systems where data privacy is crucial but collaboration across institutions is needed to enhance diagnostic accuracy. The authors in [189] also introduced a Random Forest with Self-Paced Bootstrap Learning for lung cancer prognosis, which can be adapted for kidney cancer. This framework utilizes self-paced learning to enhance the accuracy of cancer diagnosis and classification, proving that SPL can effectively handle complex datasets and improve predictive performance.
The method developed by Huang and Liang, which integrates self-paced learning for gene expression analysis, could be similarly beneficial for kidney cancer diagnosis [190] Ref4. By allowing for the gradual inclusion of challenging samples into the training set, this approach aids in identifying new, reliable biomarkers for kidney cancer, facilitating targeted therapy and personalized medicine. Shen et al. (2019) demonstrated the utility of combining deep active learning with self-paced learning for breast mass detection from mammograms [191]. This methodology can be translated into kidney cancer diagnosis, where imaging plays a crucial role, helping radiologists to detect and categorize renal masses more effectively.
In practical applications, the deep active self-paced learning model by Wang et al. [192] for accurate pulmonary nodule segmentation exemplifies how SPL can refine learning algorithms to better segment and diagnose lung nodules, a methodology that can similarly benefit kidney tumor segmentation. This approach emphasizes the active learner’s role in selecting challenging or uncertain samples, thereby optimizing the learning process and outcomes in medical image analysis. The overarching potential of SPL in medical imaging is further supported by Li et al. [193], who developed a self-paced convolutional neural network for aiding detection in medical imaging analysis. This approach could be instrumental in refining training samples for kidney cancer imaging, improving the precision of computer-aided detection systems.
Wang et al. [194] explore deep active self-paced learning for biomedical imaging, particularly focusing on pulmonary nodule segmentation in lung cancer. This method, which adjusts training focus from simple to complex samples, could similarly enhance kidney cancer diagnosis by improving model accuracy on atypical tumor presentations. Asare et al. [195] discuss a self-training scheme that incorporates self-paced learning for skin lesion classification. This approach gradually introduces harder cases into the training set, which could help develop robust diagnostic models for kidney cancer, adapting to the disease’s varied manifestations. Lastly, [196] focuses on balancing the training process between easy and challenging cases in skin disease recognition, which could be applied to kidney cancer to enhance detection accuracy and reduce biases towards common presentations.
VII-D Blockchain Technology
Blockchain technology represents a decentralized ledger that chronicles an ever-expanding list of entries known as blocks, safeguarded through cryptographic means [197]. Within the realm of cancer diagnosis, TL can be employed to educate a DL model using a vast collection of medical imagery, including CT scans, MRIs, or X-rays, for the identification of malignant tumors. After initial training, this model can undergo refinement on a narrower dataset of cancer patients to enhance its diagnostic accuracy for a particular cancer variant. A significant hurdle in deploying DL models for cancer diagnostics is the requisite for accessing and analyzing substantial volumes of sensitive patient information, all while safeguarding patient confidentiality and data integrity. Blockchain technology emerges as a pivotal solution in this context. It enables the secure, immutable storage of medical data, thus ensuring transparency and accountability for patients, medical practitioners, and researchers alike. For example, it allows patients to exercise control over their medical records, offering permissions to specific researchers or doctors, while enabling physicians to verify the data’s authenticity and precision. Moreover, blockchain facilitates the exchange and cooperative use of medical data among various healthcare entities, potentially leading to more precise diagnoses and enhanced patient care outcomes [37]. Fig.7 show the blockchain learning for KC.
Figure 7: Blockchain learning for KC
VII-E Explainable AI
The potential of explainable AI (XAI) in the diagnosis and segmentation of KC presents a promising frontier in medical imaging and oncology [198]. The challenge of interpreting complex diagnostic images, such as kidney-ureter-bladder (KUB) X-ray images, is well-documented. Ahmed et al. [199] address this by applying a TL model with a pre-trained VGG16 architecture, enhanced with XAI techniques like Layer-Wise Relevance Propagation (LRP), achieving a remarkable 97.41% accuracy in classifying KUB X-ray images as kidney stones or normal cases. This incorporation of XAI not only boosts the model’s performance but also its transparency, making the decision-making process more understandable to medical professionals. Similarly, Wani et al. [200] introduce ”DeepXplainer”, a hybrid DL model for lung cancer detection that employs SHapley Additive exPlanations (SHAP) for explicability. Although focused on lung cancer, the principles outlined could be adapted for KC diagnosis, highlighting the versatility of XAI in enhancing trust and comprehension of AI predictions in healthcare.
In kidney disease management, Moreno et al. [201] developed an explainable AI model for early detection of Chronic Kidney Disease (CKD), demonstrating how XAI can elucidate the contributions of clinical features like hemoglobin and hypertension in diagnosis. This approach can be similarly applied to KC, where understanding the influence of specific radiological features on diagnostic outcomes can significantly enhance early detection and treatment strategies. Arumugham et al. [202] and Azad et al. [203] both highlight the use of DL models combined with XAI techniques such as LIME and GRAD-CAM to improve the interpretability of predictions in kidney disease and cancer detection from various imaging modalities. These methods not only provide a heatmap or visual explanation of the AI’s focus during analysis but also reassure clinicians about the AI’s reliability, fostering a greater trust in AI-driven diagnostics. Additionally, Takahashi et al. [204] focus on the use of CNNs for detecting KC from CT images, where they employ Grad-CAM for visual explanations. This technique can be particularly useful in KC segmentation, as it highlights areas of interest that lead to a diagnosis, aiding radiologists in understanding and verifying the AI-generated outputs.
In the context of KC prediction, Lazebnik et al. [205] predict acute kidney injury (AKI) post-partial nephrectomy using machine learning algorithms with explainability components, achieving 75% accuracy. This approach underscores the critical role of preoperative parameters in predicting postoperative complications, with XAI offering a pathway to understand and improve these predictive models. In the same direction, Han et al. (2021) tackle the prediction of renal cell carcinoma (RCC) using deep convolutional TL and SHAP for model interpretability, achieving significant accuracies and providing valuable insights into model decisions [206]. This approach not only facilitates the understanding of how certain features affect the model’s prediction but also improves the clinical utility of the models by providing insights that are actionable in clinical practice.
Lastly, the research by Javaheri et al. [207] on differentiating between renal oncocytoma and chromophobe renal cell carcinoma using explainable AI frameworks emphasizes how XAI can surpass traditional CNN models in both accuracy and explainability. This is crucial for KC diagnosis, where distinguishing between different types of renal tumors can significantly impact treatment decisions.
Table VII provides a summary of research on ML models enhanced with XAI in KC diagnosis and related analysis. The incorporation of XAI into kidney cancer diagnosis models, as highlighted in the table, significantly enhances the transparency and interpretability of machine learning predictions, which is critical in medical decision-making. Key findings include the successful integration of XAI techniques like SHAP, LRP, and LIME that improve the fairness, understandability, and reliability of model outputs. These XAI-enhanced models facilitate a deeper understanding of how specific features influence diagnostic predictions, thus fostering trust among clinicians and aiding in the effective communication of diagnostic reasoning. For instance, models have achieved high diagnostic accuracies (up to 99.2%) while providing actionable insights into feature relevance, such as the significance of hemoglobin levels in CKD prognosis. This integration addresses a crucial gap in traditional AI approaches by making advanced AI diagnostics more accessible and actionable in clinical settings, particularly enhancing the early detection and treatment planning for kidney-related diseases.
VII-F Adversarial Training
The burgeoning field of adversarial training in DL offers a compelling avenue for enhancing the diagnosis, prediction, and segmentation of KC. This approach leverages the strengths of generative adversarial networks (GANs) and adversarial training methods to address some of the inherent challenges in medical imaging, such as data scarcity, variability in tumor appearance, and the robustness of diagnostic models against adversarial examples. In this respect, Zheng et al. [208] introduce a multi-scale adversarial learning approach for the segmentation of kidney and kidney tumor areas. Their method, which incorporates difficult region-supervised loss and adversarial learning at multiple scales, shows promise in accurately delineating complex tumor shapes and sizes, which are often challenging due to their diverse appearances and fuzzy boundaries. Such an approach is crucial for precise radiotherapy planning and monitoring disease progression, ensuring that treatments are targeted and effective. The MB-FSGAN model discussed by Ruan et al. [209] integrates multi-scale feature extraction and generative adversarial learning to enhance the segmentation and quantification of kidney tumors. This method not only segments but also quantifies tumor indices, providing critical morphometric details that are essential for personalized treatment planning. The adversarial component helps in refining the segmentation accuracy, proving especially useful in clinical settings where precision is paramount. Shan et al. [210] developed SegTGAN, a GAN-based model for autonomous kidney segmentation, which outperformed traditional networks like U-Net and FCN on the Kits19 dataset. The incorporation of multi-scale feature extraction and densely connected blocks in a GAN framework suggests that adversarial training can significantly improve the accuracy of segmentation models, a critical step for effective diagnosis and treatment planning in KC. In the study by Gadermayr et al. [148], the application of adversarial models for image-to-image translation demonstrates a significant advancement in the field of digital pathology segmentation. This approach effectively addresses the variability in staining—a common challenge in pathology—without requiring extensive manual annotations for each characteristic. By enabling stain-independent supervised segmentation and combining it with unsupervised techniques, the study showcases how adversarial training can adaptively manage inherent variabilities in pathological images, leading to improved segmentation outcomes that are crucial for accurate cancer diagnosis and treatment planning. Moving on, Zeng et al. [211] utilize an unsupervised domain adaptation technique to address the segmentation of kidney structures from computed tomography urography images. This method tackles the domain shift problem effectively, demonstrating that adversarial training can bridge the gap between different imaging modalities. By enabling a model to learn deeply from urographic images without direct supervision, this approach provides a way to enhance the 3D reconstruction and visualization of kidneys, which is vital for surgical planning and outcome analysis in KC treatment.
Besides, Uhm et al. [212] propose a unified framework that uses adversarial learning to generate missing CT phases, which are then used to classify KC subtypes. This approach demonstrates how adversarial training can not only enhance data completeness but also improve diagnostic accuracies by enabling the use of multi-phase CT scans that provide comprehensive views of the kidney, essential for accurate diagnosis and staging of cancer. On the other hand, the research by Brinda et al. [213] illustrates the application of Conditional Variational Generative Adversarial Networks (CVGANs) in the diagnosis of chronic kidney disease, which often precedes or accompanies KC. By optimizing the selection of diagnostic features through a metaheuristic algorithm and training the model to classify disease presence more accurately, this study highlights the potential of adversarial networks to refine the prediction capabilities of diagnostic models. The high precision and recall achieved underscore the effectiveness of GANs in enhancing the reliability of computer-aided diagnostic systems. Similarly, the study by Sughasiny et al. [214] introduces the Deep Adversarial Belief Network (DABN), which leverages adversarial training to enhance the classification of renal cell carcinoma (RCC). This approach is significant because RCC is often detected in its late stages, making effective treatment more challenging. The DABN model utilizes adversarial inputs to fine-tune the network’s ability to discern between healthy and cancerous tissue, even when the differences are subtle. This method not only improves accuracy but also addresses the problem of tuning hyperparameters and managing outliers that often affect DL models. By comparing performance with existing methods like E-CNN and FPSO CNN, the DABN demonstrates a higher accuracy, illustrating the potential of adversarial models in refining diagnostic algorithms for early and more precise detection of KC. Moving forward, the work by Liao et al. [215] showcases how DL, specifically through the use of ACWGAN-GP and MobileNetV2, can effectively tackle the problem of imbalanced and insufficient data samples in the classification of chronic kidney disease stages. Adversarial networks, by generating additional training data, can help in creating more balanced datasets that enhance the robustness and generalization ability of diagnostic models. This method could similarly be adapted for KC, where stages or types of cancer could be more accurately classified with augmented training sets.
TABLE VII: Summary of research on ML models enhanced with XAI in KC diagnosis and related analysis.
Ref. ML Model Used Dataset Contribution Description Best Performance Value Limitation
[198] NN-CBR twin system Population health data Implemented a neural network complemented by CBR for explainable predictions of CKD risk, aimed at improving healthcare predictions. Accuracy of 95% Neural networks remain opaque without effective implementation of explainable systems.
[199] VGG16 with XAI (LRP) KUB X-ray images Developed a system for classifying kidney stones using transfer learning and XAI to enhance transparency and decision-making. Accuracy of 97.41% Lacks fairness and explainability without XAI integration.
[200] ConvXGB with SHAP for XAI Survey Lung Cancer dataset Introduced ”DeepXplainer,” a hybrid deep learning model for lung cancer detection and explanation, integrating CNNs and XGBoost with SHAP for explainability. Accuracy of 97.43%, Sensitivity of 98.71%, F1-score of 98.08% The complexity of integrating multiple AI technologies for explainability.
[201] Extreme Gradient Boosting with XAI Clinical features dataset Developed an explainable CKD prediction model using a minimal set of clinical features, optimized for early diagnosis. Accuracy of 99.2% with cross-validation Reliance on a limited number of clinical features which may not capture all nuances of CKD.
[202] DNN with LIME for XAI Healthcare diagnostic data Built a DNN model to predict early-stage CKD with explainability provided by LIME, focusing on interpretability in healthcare applications. Accuracy of 98.75%, ROC_AUC of 98.86% The potential black-box nature of DNN without effective explainability.
[203] Deep learning with GRAD-CAM Robotic nephrectomy videos Utilized deep learning to detect tumors in laparoscopic images, enhancing tumor localization and classification. mAP of 0.974, classification accuracy of 0.84 Specific to images from surgical procedures, may not generalize to other forms of imaging.
[205] Random Forest with Boolean decision trees Data from patients undergoing open PN Predicted AKI post-partial nephrectomy using a Random Forest model optimized for pre-operative parameters. Accuracy of 75% Limited to predictions based on pre-operative data only.
[206] Deep convolutional transfer learning with SHAP Hunan Cancer Hospital and COIL-100 datasets Proposed an RCC prediction model incorporating SHAP for explainability and DCA for evaluating clinical utility. Accuracy of 73.87% on RCC, 99.81% on COIL-100 Lower performance in real-world clinical datasets compared to public datasets.
[207] Explainable AI framework H&E-stained images of ChRCC and RO Developed an explainable framework to differentiate between ChRCC and RO, enhancing diagnostic accuracy and reducing pathologist workload. Diagnostic accuracy of 88.2% The challenge of differentiating similar cancer types based on microscopic features.
[204] CNN with Grad-CAM Abdominal CT images Explored the use of CNNs and visual explanations via Grad-CAM to detect kidney cancer from CT images. Not specified Challenges in handling the diversity of abdominal CT images.
[216] XGBoost with SHAP for XAI Expression data of skin samples Applied XAI to identify significant genes for SCC diagnosis, providing a basis for potential biomarkers. Not specified Specific to non-melanoma skin cancers and dependent on data quality.
[217] Extreme Gradient Boosting with XAI CKD indicators dataset Developed an explainable CKD prediction model, balancing accuracy and explainability using ensemble trees and selected features. Accuracy of 99.2%, cross-validation accuracy of 97.5% Focus on a small set of features, which may not apply universally.
Lastly, in a parallel approach, Hu et al. [218] focus on the issue of image artifacts in prostate cancer diagnosis, which shares similarities with KC imaging challenges. They introduce a Targeted Adversarial Training with Proprietary Adversarial Samples (TPAS) strategy to combat the misleading effects of rectal artifacts in MRI scans. This method specifically generates adversarial samples that mimic these artifacts, training the system to recognize and disregard such misleading information. This adversarial approach enhances the model’s robustness, ensuring that the diagnostic accuracy is not compromised by common imaging anomalies. This strategy underscores the utility of adversarial training in improving the reliability of cancer diagnostics by making the models resilient to specific challenges inherent in medical imaging.
Figure 8: Federated learning for KC.
Table VIII summarizes the relevant research on adversarial training in kidney disease and tumor analysis. It is has been shown that the use of adversarial learning strategies allows for better handling of the inherent variability in tumor appearance, resulting in more precise segmentation outcomes, as demonstrated in studies using models like MSALDS-UNet [208] for fine segmentation tasks. Moreover, adversarial training aids in refining models to detect and segment challenging and less distinct regions, such as fuzzy tumor boundaries, thereby supporting more accurate clinical assessments and potential treatment planning enhancements. These advancements highlight the transformative potential of adversarial training techniques in dealing with the complex imaging characteristics of kidney tumors.
TABLE VIII: Summary of research on adversarial training in KC diagnosis.
Ref. ML Model Used Dataset Contribution Description Best Performance Value Limitation
[215] ACWGAN-GP, MobileNetV2 Renal ultrasound images Developed an auxiliary diagnosis system for CKD stage classification using deep learning. Accuracy up to 90.1% Limited to four stages of CKD.
[208] MSALDS-UNet KiTS21 Proposed a cascade-based approach with multi-scale adversarial learning for kidney and tumor segmentation. Not specified Handling diverse tumor shapes and locations.
[209] MB-FSGAN CT scans of 113 kidney tumor patients Introduced a GAN for simultaneous segmentation and quantification of kidney tumors. Pixel accuracy of 95.7% Manual segmentation variability and time consumption.
[210] SegTGAN Kits19 dataset Developed a GAN-based technique for autonomous kidney segmentation. DSC of 92.28%, ACC of 97.28% Quantitative analysis complexity due to data scarcity.
[212] Fully 3D convolutional networks In-house and external datasets Created a framework for kidney cancer diagnosis with incomplete multi-phase CT. Higher performance than using incomplete data Relies on availability of complete multi-phase CT.
[214] Deep Adversarial Belief Network (DABN) Not specified Utilized Grasshopper Optimization Algorithms for kidney cancer detection. Outperforms existing methods (up to 98.25% accuracy) Hyper-parameter tuning and outlier effects.
[218] Targeted adversarial training with Proprietary Adversarial Samples Not specified Improved prostate cancer classification by mitigating rectal artifact interference. Substantial gains over ordinary training Specific to prostate MRI and adversarial samples.
[148] Image-to-image translation models Patches showing kidney histology Developed unsupervised segmentation approaches for digital pathology. Significant improvements in unsupervised segmentation Requires handling of high image variability.
[211] Unsupervised domain adaptation with 2D networks Clinical urography data Proposed unsupervised kidney segmentation from urographic images. Comparable or better results than supervised methods Adaptation issues across clinical unseen data.
[213] Conditional Variational Generative Adversarial Networks Benchmark CKD dataset Developed a feature selection-based classifier for CKD diagnosis using deep learning. Accuracy of 99.2%, Recall of 98.6% Pre-processing to remove data with absence of values.
VII-G Tranformers
The integration of transformer architectures, such as bidirectional encoder representations from Transformer (BERT) and connectionist temporal classification (CTC) among others [219], into the realm of KC diagnosis and segmentation represents a cutting-edge approach that holds significant promise for advancing medical imaging analysis. Drawing insights from several studies, such as those referenced, the potential applications of transformers in this domain are evident. One of the primary advantages of transformers lies in their ability to capture long-range dependencies within the data, making them particularly suitable for tasks involving complex relationships and patterns, as often encountered in medical imaging. The referenced studies showcase various transformer-based models tailored specifically for KC diagnosis and segmentation.
For instance, the utilization of convolution-and-transformer networks (COTRNet) proposed by Shen et al. [220] offers an end-to-end solution for kidney tumor segmentation, demonstrating the effectiveness of combining convolutional and transformer-based architectures in this context. Similarly, Islam et al. [54] leverage vision transformers for the automatic detection of kidney cysts, stones, and tumors from CT radiography, highlighting the versatility of transformer models across different types of kidney abnormalities.
Furthermore, the study by Wang et al. [221] showcases the synergy between transformers and convolutional neural networks (CNNs) for renal parenchymal tumors diagnosis. By stacking transformers and CNNs, the model efficiently captures both long-range dependencies and low-level spatial details, enhancing its performance in identifying and diagnosing kidney lesions. Moreover, transformer-based methods, such as TA-UNet3+ proposed by Hu [222], introduce innovative attention mechanisms and encoder-decoder architectures to improve kidney tumor segmentation accuracy. Similarly, Qian et al. [223] present a hybrid network combining nnU-Net and Swin Transformer, emphasizing the critical role of precise delineation and localization in KC diagnosis and treatment planning. Additionally, transformers exhibit promise in histopathological image analysis for subtype classification of renal cell carcinoma (RCC). Gao et al. [224] demonstrate the effectiveness of instance-based vision transformers in subtyping papillary RCC, providing valuable insights for histopathological diagnosis and personalized treatment strategies.
VII-H Quantum Computing
Quantum computing represents a nascent and burgeoning domain poised to transform our approach to computational challenges, including those within ML, with the potential to revolutionize various aspects of knowledge discovery [225]. Despite being at an early developmental stage, quantum computing harbors the promise of significantly enhancing the efficacy of TL models in the diagnosis of diseases like KC [226]. Employing quantum computing in TL for KC diagnosis entails harnessing quantum algorithms to tackle tasks that would otherwise be impractical with classical computers, notably optimization and linear algebra computations. These algorithms offer avenues to bolster TL model performance by curtailing computational overheads and enhancing model accuracy [227]. Additionally, quantum computing holds promise in mitigating classical computing constraints encountered in medical imaging, such as the demand for substantial computational resources and memory for processing voluminous datasets [228, 229]. By leveraging quantum computing, the duration required for training and inference can be slashed, fostering swifter diagnoses and improved patient outcomes [230, 231].
While quantum computing holds immense promise for revolutionizing cancer diagnosis using transfer and DL techniques, significant advancements in quantum algorithms, hardware capabilities, and data availability are necessary to realize its full potential in improving the diagnosis and treatment of KC and other types of cancer [232]. Collaborative efforts between quantum physicists, computer scientists, and medical researchers are essential to overcome these challenges and pave the way for the future of quantum-enhanced healthcare technologies [233].
VII-I Large Language Models and Generative Chatbots
Large language models (LLMs) and generative chatbots hold significant potential for transforming KC diagnosis in several ways [234, 235]. By integrating vast medical libraries and research data, LLMs can provide healthcare professionals with up-to-date information, diagnostic suggestions, and treatment options based on the latest studies and clinical guidelines [236, 237]. Moving on, generative chatbots can interpret patients’ descriptions of their symptoms in natural language, aiding in the early detection of KC by flagging potential warning signs and suggesting further diagnostic tests [238, 239, 240]. Typically, these models can deliver personalized information to patients about KC, including explanations of their condition, possible treatment pathways, and what to expect during each treatment option, improving patient understanding and engagement in their care plan [241, 242, 243].
Besides, LLMs can analyze vast amounts of data from EHRs to identify patterns or risk factors associated with KC, potentially uncovering new insights into its diagnosis and progression [244, 245]. With training, LLMs can assist radiologists by providing a second opinion on imaging studies, such as CT scans and MRIs, identifying areas that may warrant a closer look for signs of KC [246, 247, 248]. Additionally, chatbots can handle routine inquiries and patient triaging, allowing medical staff to focus on more complex tasks and patient care, thereby increasing the efficiency of clinical workflows [249, 250]. LLMs can integrate and analyze information from various sources, including genomic data, radiographic images, and clinical notes, offering a holistic view of a patient’s health status and aiding in the accurate diagnosis of KC [251, 252]. Moreover, generative chatbots can facilitate remote monitoring of patient’s symptoms and treatment responses, providing timely advice and enabling telehealth consultations with specialists [253, 254].
VIII Conclusion
The application of TL and adaptive domain techniques presents a promising avenue for enhancing the precision of KC diagnoses. TL enables the adaptation of models trained on extensive datasets to new, smaller datasets where labeled data may be scarce. Meanwhile, adaptive domain techniques allow models to adjust to shifts in data distribution over time, a critical factor for maintaining accuracy in the dynamic landscape of medical diagnostics. By employing TL, models pre-trained on comprehensive medical image datasets can undergo fine-tuning with smaller sets of kidney images for more precise KC diagnosis. This approach leverages the pre-training phase’s acquired relevant features, proving advantageous when faced with limited labeled data availability. The adaptive domain further refines model accuracy as data distribution evolves, a key aspect in KC diagnostics where new data continuously emerges. Ensuring the model’s adaptability to these changes is essential for sustaining diagnostic precision.
Our research delves into the utilization of TL in identifying KC, highlighting its potential to mitigate computational demands compared to conventional DL models, circumvent annotated data shortages, and handle dataset variability across training and testing phases. The paper’s organizational section lays out a detailed discourse on various concepts and classifications. It deep-dives into the literature surrounding TL-driven KC detection, offering an extensive classification of contributions to the field. Subsequently, it proposes several general solutions to the challenges encountered in KC detection, applicable to both traditional and TL-based models. Drawing from the reviewed literature, it advocates for future methodological advancements aimed at precise and accurate KC detection, underscoring the importance of harnessing TL’s full spectrum of capabilities.
Despite the successes of TL and DA in KC detection, numerous challenges warrant further exploration. One such challenge is the homogeneous nature of many DA models, which assume consistent feature distances between source and target domains—an assumption not always valid. Some studies have ventured into knowledge transfer without this prerequisite, utilizing datasets to enhance task support. Moreover, as highlighted, the future is likely to see heightened focus on heterogeneous adaptive domain approaches until their optimization is achieved, thereby advancing KC detection systems.
Data availability
Data will be made available on request.
Conflict of Interest
The authors declare no conflicts of interest.