### v2

Abstract
Incurable diseases continue to pose significant challenges to global healthcare systems, with their prevalence being influenced by lifestyle choices, economic conditions, social factors, and genetic predispositions. Among these, kidney disease remains a critical health issue, affecting individuals worldwide and necessitating continuous research to enhance early diagnosis and treatment strategies. In recent years, deep learning (DL) has gained prominence in medical imaging and diagnostics, offering significant advancements in automatic kidney cancer (KC) detection. However, the effectiveness of DL models largely depends on the availability of high-quality medical datasets, which are often scarce, expensive, and domain-specific. Additionally, these models require substantial computational resources and storage capacity, limiting their applicability in real-world clinical settings. To address these challenges, transfer learning (TL) has emerged as a promising solution, enabling the adaptation of pre-trained models from different domains to improve KC diagnosis. This paper presents the first comprehensive survey of DL-based TL frameworks specifically applied to KC detection. By systematically reviewing existing approaches, this study identifies key methodologies, explores their advantages and limitations, and critically analyzes their effectiveness in clinical applications. Furthermore, the paper highlights current challenges in implementing TL for medical imaging and discusses emerging trends that could shape future research in this field. Ultimately, this review underscores the transformative role of TL in precision medicine, particularly in oncology, by enhancing diagnostic accuracy, reducing computational costs, and facilitating the deployment of AI-driven healthcare solutions. The findings of this study provide valuable insights for researchers and practitioners, paving the way for further advancements in KC diagnostics and personalized treatment approaches.
keywords:
Medical image, Kidney cancer diagnosis, Deep learning, Transfer learning, Domain adaptation, Fine-tuning.
1 Introduction
The distinction between tumor subtypes and pathological grades is a critical aspect of cancer diagnosis and treatment planning. Tumor subtypes classify the tumor based on its cellular and molecular characteristics. For instance, in renal cell carcinomas (RCC), the major subtypes— clear cell renal cell carcinoma (CCRCC), papillary renal cell carcinoma (PRCC), and chromophobe renal cell carcinoma (CRCC)—differ in their cellular appearance, growth patterns, and genetic alterations. These subtypes have distinct biological behaviors, prognoses, and responses to treatment. RCC is a cancer that impacts kidney function, which plays a vital role in regulating the body’s fluid balance by filtering out toxins and excreting waste products. Disorders such as kidney cancer (KC) can impair this function [1]. As shown in Figure 1, the estimated incidence and mortality rates for KC in 2022, for both genders globally. In terms of pathological grading, KC is classified on a scale of 1 to 4, with 1 being the least aggressive and 4 being the most aggressive. This grading is determined by analyzing cancer cells under a microscope to assess the degree of differentiation and anaplasia. Grade 1 (PRCC) tumors are well-differentiated, closely resembling normal kidney cells and grow slowly [2]. Grade 2 (CRCC) tumors are moderately differentiated with some resemblance to normal kidney cells, growing more slowly than grades 3 and 4. Grade 3 tumors (CCRCC) consist of poorly differentiated cells that bear little resemblance to normal kidney cells and tend to grow and spread rapidly [3]. Grade 4 tumors ( sarcomatoid renal cell carcinoma (SRCC)) are highly aggressive and rare, growing and spreading rapidly. Pathological grading, which reflects how abnormal the tumor cells are compared to normal tissue, helps predict the tumor’s aggressiveness and guides therapeutic decisions. While subtypes categorize the tumor based on its origin and intrinsic characteristics, grading evaluates the degree of malignancy. Understanding both the subtypes and grades is crucial for determining prognosis and selecting the most effective treatment strategy. We have also referenced the WHO pathological classification and grading system for renal tumors [4], providing a reliable and standardized framework for understanding these concepts. These revisions enhance the scientific rigor and clinical relevance of our paper, making it more valuable for researchers and healthcare professionals alike.
Acronyms and Abbreviations
AI
artificial intelligence
BERT
bidirectional encoder representations from Transformer
CNN
convolutional neural network
CT
computerized tomography
CTC
connectionist temporal classification
CRCC
chromophobe renal cell carcinoma
CCRCC
clear cell renal cell carcinoma
CAD
computer-aided detection
DA
domain adaptation
DG
domain generalization
DL
deep learning
FSL
few-shot learning
FL
federated learning
KC
kidney cancer
KiTS19
kidney tumor segmentation 19
LRP
layer-wise relevance propagation
LLM
large language model
ML
machine learning
MRI
magnetic resonance imaging
PRCC
papillary renal cell carcinoma
RCC
renal cell carcinomas
RDAN
reciprocal domain adaptation network
SVM
support vector machine
SRCC
sarcomatoid renal cell carcinoma
SSL
self-supervised learning
TL
transfer learning
TCGA
the cancer genome atlas
US
ultrasound
UL
unsupervised learning
XAI
explainable AI
Figure 1: The estimated number of KC incidence and mortality in the world.
Moreover, KC can be identified by employing a range of imaging methods such as magnetic resonance imaging (MRI), X-rays, and ultrasound (US) imaging [5]. These methods are capable of generating comprehensive visuals of the kidneys, enabling physicians to spot any irregularities or tumors [6, 7]. Nonetheless, despite the utility of medical imagery in disease diagnosis, its reliability is not absolute in certain scenarios. Particularly in cancer cases, more precise techniques may be required to aid physicians in detecting malignant cells and assessing their metastasis to other body regions [8]. Moreover, interpreting whether an image truly signifies the presence of cancer can be challenging. Medical imagery is often rich in detail and data, complicating its analysis. In addition, the overwhelming volume of images and visual assessments may contribute to medical inaccuracies. The burden of analyzing numerous images can diminish a radiologist’s focus and their capacity to notice crucial nuances. Such circumstances can lead to misdiagnosis, diagnosis delays, or even inappropriate treatments [8].
In recent years, deep learning (DL)-based early detection methods have recorded significant achievements in identifying various terminal illnesses. These methods focus on accurately extracting diverse features that assist medical professionals in the diagnostic process through the utilization of the network’s multiple layers [9, 10]. Attaining the necessary diagnostic precision for incurable diseases affecting the kidneys presents challenges due to: (i) the insufficiency of data resources for the training phase [11, 12]; (ii) the time-consuming, expensive, and exhaustive nature of data collection [13, 14]; (iii) the requirement for data annotation in the training phase; and (iv) the impracticality of manually labeling large datasets by experts, which is often unfeasible [15, 16].
While many DL algorithms, such as the convolutional neural network (CNN) [17, 18, 19, 20], are proficient at recognizing complex patterns, they encounter significant challenges, particularly during the transition from training to testing phases, which can impede their effectiveness. Additionally, various studies [21] have pointed out that the majority of datasets utilized for diagnosing either early or advanced cancer stages are often inadequate for the diagnostic process. This inadequacy is due to the intricate structure of certain organs, like the kidneys, and the variations in image depth, making it exceedingly difficult to accurately identify the type or grade of cancer [22, 23, 24]. Furthermore, these algorithms heavily rely on the assumption that the training and testing data are extracted from identical feature spaces and distributions. A deviation between these can necessitate a comprehensive model reconfiguration, which is not only costly but also labor-intensive. As [25] notes, supervised learning algorithms can yield remarkable outcomes with extensive volumes of annotated data. Yet, they generally falter with the complex variations present in medical images, where labeled data are scarce, and backgrounds are complex. Conversely, diagnostic systems powered by artificial intelligence (AI) critically depend on having access to well-distributed and annotated datasets for training, a requirement that is often hard to fulfill [26]. This underscores the need for enhancements in AI-based diagnostic systems for incurable diseases, aiming to simplify them to improve diagnostic efficiency, timing, and precision [27].
To meet these goals, efforts are being directed towards leveraging transfer learning (TL) in many research fields [28, 29, 30, 31, 32]. This process fundamentally involves repurposing a network that has been previously trained on a specific dataset for a distinct task. This entails employing established models such as LeNet, AlexNet, VGGNet, ResNet, GoogLeNet, DenseNet, XceptionNet, and SqueezeNet [29] for tasks on smaller datasets, thereby facilitating the omission of labels. This concept draws inspiration from the human ability to apply knowledge and experiences from past learning to foster new insights. Consequently, various databases hosting extensive image collections have been made available online, such as the CNN algorithm’s application on the ImageNet dataset, which is accessible for public experimentation with DL-based algorithms. This approach effectively mitigates data scarcity issues and circumvents the need for extensive data annotation [33]. Moreover, this algorithmic approach alleviates the lengthy durations typically required for learning processes, which can extend over days or weeks. In the realm of diagnosing terminal illnesses, TL has been categorized into two methodologies: fine-tuning and feature extraction, both of which have shown promising outcomes, notably in KC detection. Numerous studies within this domain, for instance, [34], exemplify its application. TL is particularly beneficial for cancer detection and diagnosis, including KC [35, 36]. Utilizing a pre-trained DL model, like a CNN, fine-tuned on kidney imaging scans (e.g., computerized tomography (CT) or MRI), can enhance its efficacy in identifying and diagnosing kidney tumors (Figure  2). A method for fine-tuning a model for KC detection might involve using a dataset comprising both healthy and cancerous kidney scans, adjusting the model’s final layers’ weights to better recognize cancerous areas. Additionally, the model could be optimized by either freezing specific layers while training the last few or starting a new one with the pre-trained weights. An alternative TL strategy in KC detection could leverage a model pre-tuned for a similar task, such as lung cancer detection, and further fine-tune it on kidney scans, beneficial when large datasets for KC detection are unavailable. TL proves invaluable in medical imaging, significantly when large, labeled datasets are scarce, enhancing the efficiency and accuracy of kidney tumor identification and diagnosis.
Figure 2: Principle of Kidney classification based on TL.
This study is the first comprehensive review based on DL and TL for the diagnosis of KC. While many studies have explored the application of DL in medical imaging, few have systematically addressed the integration of TL specifically for KC detection. Previous reviews either focus on DL alone or do not consider TL frameworks, which are particularly important for overcoming challenges like dataset scarcity and limited computational resources in real-world clinical settings. This paper fills that gap by presenting a thorough survey of DL-based TL frameworks applied to KC detection, systematically analyzing existing approaches and identifying their advantages and limitations. It goes further by addressing challenges such as the availability of high-quality medical datasets, the need for extensive computational resources, and the difficulties in adapting DL models for practical use in clinical environments. Additionally, the paper highlights emerging trends in TL, which could play a crucial role in enhancing diagnostic accuracy and facilitating AI-driven healthcare solutions. Thus, this review not only addresses a significant gap in the literature but also paves the way for future advancements in KC diagnostics, offering valuable insights for researchers and healthcare practitioners.
1.1 Our Contributions
While numerous review articles have examined the applications of TL algorithms in medical imaging and disease detection—such as [34, 37, 38]—to the best of the authors’ knowledge, no comprehensive survey has been conducted that specifically addresses the role of TL in the context of kidney cancer (KC). This paper aims to fill that gap by offering a novel and unified examination of how TL and domain adaptation ( domain adaptation (DA)) techniques contribute to the detection and management of KC.
Our work provides a structured and in-depth review of TL’s applications across different types and stages of kidney cancer, integrating classification strategies and key conceptual frameworks. This unified perspective aims to demystify the role of TL in KC detection, with a particular focus on enhancements through fine-tuning and cross-domain DA. Furthermore, this review underscores the importance of TL in multi-source data fusion to improve diagnostic accuracy, and it highlights open challenges and future research directions, advocating for the development of more generalizable and computationally efficient TL models for KC detection. The unique contributions of this review are as follows:
  • Exploring various facets of TL and DA that have driven their evolution in medical imaging.
  • Providing a detailed taxonomy of TL and DA frameworks, accompanied by an analysis of their foundational concepts.
  • Conducting a comprehensive literature review on the role of TL and DA in KC detection, based on consistent evaluation criteria. Demonstrating how TL and DA approaches can enhance integrated data pipelines for improved KC diagnosis.
  • Highlighting unresolved challenges in the application of TL and DA to cancer detection, including: (1) model generalizability and precision, (2) evaluation metric selection, (3) the risk of transferring task-irrelevant knowledge, among others.
  • Proposing future research directions aimed at lowering computational complexity and resource demands in TL-based frameworks.
Establishing this work as the first survey, to the best of the authors’ knowledge, to systematically address the integration of TL and DA within the KC detection landscape. It contributes uniquely by (1) investigating diverse applications, (2) presenting an integrated framework, (3) identifying technical and methodological challenges, and (4) offering insights for future exploration. Table. 1, displays the results of the comparison of the contribution of the proposed review with other surveys in the field of TL. This study is intended for a diverse audience that includes researchers, healthcare professionals, and technical experts working in the fields of medical imaging, oncology, and AI. It is particularly relevant for those interested in leveraging cutting-edge DL and TL techniques to improve early detection and diagnosis of KC. This audience may consist of biomedical engineers, data scientists, and oncologists who are looking to enhance diagnostic accuracy, overcome challenges in dataset limitations, and explore innovative methodologies for precision medicine. Furthermore, the study is valuable for academics and industry professionals aiming to integrate advanced AI-driven technologies into clinical practices for cancer diagnosis and treatment planning.
Table 1: A summary of a comparison of the proposed study’s contribution with other studies in the field of TL is shown. A check mark (✓) signifies that the area has been covered, while a cross mark () indicates that it has not been addressed.
Review Year Description TLB TLS DA TLA TLD TLAD TLL KCH FD
[39] 2018 TL for mental health
[40] 2018 TL for cancer detection
[41] 2019 TL for breast cancer
[42] 2020 TL for medical image
[43] 2021 TL for EEG signal analysis
[44] 2021 TL for MRI brain
[45] 2021 TL for Alzheimer’s disease
[46] 2022 TL for neuroimaging analysis
[47] 2022 TL for KC
[48] 2023 TL for COVID-19 medical image
[49] 2023 TL for Diabetic Retinopathy
Our 2024 TL for KC
Abbreviations: TL background (TLB), TL segmentation (TLS), TL applications (TLA), TL datasets (TLD), TL advantages (TLAD), TL limitations (TLL), Key challenges (KCH), Futur direction (FD)
2 Methodology
2.1 Search strategy and paper selection
This survey on deep TL for KC diagnosis provides a comprehensive exploration of existing DL-based TL frameworks and their application in the medical field, specifically targeting KC diagnosis. The methodological approach taken in this review encompasses several systematic steps to ensure a thorough analysis and presentation of the current state of research in this area.
The literature search was conducted using several academic databases including PubMed, IEEE Xplore, Scopus, Springer Nature, Elsevier, Wiley, Taylor & Francis, ACM, and Google Scholar. The search was designed to identify relevant articles published in the last five years, ensuring the timeliness and relevance of the data collected. The keywords used for the search included combinations of “TL”, “deep TL”, “DA”, “KC diagnosis”, and “medical imaging”. The search strategy was adapted to the syntax and capabilities of each database to maximize the retrieval of pertinent articles. The main inclusion and exclusion criteria, are:
  • Inclusion criteria:
    • Studies that specifically apply DL and TL techniques.
    • Articles focusing on the diagnosis of KC using medical imaging.
    • Papers published within the last five years to ensure current relevance.
    • Studies providing clear methodological details and quantitative results.
  • Exclusion criteria:
    • Non-English publications.
    • Conference abstracts or presentations without full-text availability.
    • Studies not involving explicit use of DL or TL frameworks.
    • Papers not directly related to KC.
2.2 Bibliometric Analysis
To explore and scrutinize the scientific investigations detailed in this review, bibliometric methods were utilized. The cumulative interest in TL-based KC over time is summarized in Figure 3. This illustration provides a snapshot of the enduring interest in research on TL-based KC, highlighting a notable uptick in enthusiasm for crafting TL-based KC methodologies. Specifically, this surge in interest is represented in Figure 3(a), indicating a sharp rise in the publication count, which peaked at 66 articles in 2023. Furthermore, Figure 3(b) displays the countries contributing most significantly to research outputs in this area, with China and the United States emerging as the most prolific in TTL-based KC research advancements. Furthermore, Figure 3(c) identifies the differents domain of application of KC-oriented TL. Lastly, Figure 3(d) delineates the distribution of paper types, revealing that journal articles predominate, constituting 65.8% of all published work, with conference papers making up 19.7%.
Figure 3: Bibliometric analysis in terms of: (a) The number of published articles; (b) The most country known by active researchers; (c) The percentage of the contribution in each domain; and (d) Distribution of paper types.
Initial screening was based on titles and abstracts, with articles potentially meeting the inclusion criteria selected for full-text review. This stage filtered out irrelevant or out-of-scope studies, streamlining the focus on TL in the context of KC. Essential information was extracted from each article, including study objectives, DL models, datasets, TL techniques employed, and principal findings. This structured approach facilitated comparative analysis and synthesis.
The structure of this document is outlined below. Section 2 describes the methodology used in conducting this review, including the search strategy and bibliometric analysis. Section 3 presents a comprehensive taxonomy of TL and domain adaptation (DA) techniques applied in the context of KC, including inductive, transductive, domain generalization, self-taught, and few-shot learning, along with associated strategies such as fine-tuning and multi-task learning. Section 4 explores key applications of TL in KC, encompassing cancer detection, CT scan analysis and segmentation, genomic data analysis, multi-modal data fusion, and explainable AI. Section 5 provides an overview of the most commonly used datasets and evaluation metrics in this domain. Challenges encountered in employing TL for KC diagnosis and potential directions for future research are discussed in Sections 6 and 7, respectively. Finally, Section 8 summarizes the conclusions drawn from this comprehensive review.
3 Taxonomy of TL
TL is a method in machine learning (ML) that involves applying knowledge obtained from solving one problem to a different, yet related, problem. This part presents the mathematical framework underlying TL, beginning with the concepts of a "domain" and a "task."
A domain $\mathcal{D}$ consists of two primary elements:
  • A feature space $\mathcal{X}$, encompassing all potential inputs.
  • A marginal probability distribution $P(X)$, with $X\in\mathcal{X}$, which describes the distribution of samples across the feature space.
A task $\mathcal{T}$ is distinguished by: A label space $\mathcal{Y}$; A goal-oriented predictive function $f(\cdot)$ that associates an input $x\in\mathcal{X}$ with an output $y\in\mathcal{Y}$. This function $f$ is generally unknown and represents the target of learning algorithms to estimate.
In TL, we deal with a source domain $\mathcal{D}_{S}$ and a target domain $\mathcal{D}_{T}$, along with their corresponding tasks $\mathcal{T}_{S}$ and $\mathcal{T}_{T}$. The objective is to improve the learning of the target predictive function $f_{T}(\cdot)$ in $\mathcal{D}_{T}$ using the knowledge from $\mathcal{D}_{S}$ and $\mathcal{T}_{S}$.
In the context of TL, we engage with both a source domain $\mathcal{D}_{S}$ and a target domain $\mathcal{D}_{T}$, as well as their respective tasks $\mathcal{T}_{S}$ and $\mathcal{T}_{T}$. The aim is to enhance the learning of the target predictive function $f_{T}(\cdot)$ within $\mathcal{D}_{T}$ by leveraging the knowledge from $\mathcal{D}_{S}$ and $\mathcal{T}_{S}$.
The purpose of a TL approach is to utilize the source domain and task to improve the accuracy of the predictive function in the target domain. This involves adapting models, identifying shared feature representations, or employing other methods to make use of the similarities between the domains and tasks. Table 2, summarizes the types of TL. Table 3 provides a summary of various cutting-edge techniques that utilize TL types for KC, including the datasets and metrics employed.
Table 2: Types of TL.
TL settings Source Target Tasks
Traditional ML/DL $D_{s}=D_{T}$ $D_{s}=D_{T}$ All
Inductive TL $D_{s}\cong D_{T}$ $D_{s}\neq D_{T}$ Classification and regression
Transductive TL $D_{s}\cong D_{T}$ $D_{s}=D_{T}$ Classification and regression
Unsupervised TL $D_{s}\neq D_{T}$ $D_{s}\neq D_{T}$ Clusterings and reduction
3.1 Inductive TL
Inductive TL focuses on using the knowledge gained from the source domain and task to improve learning in a related but different target task. The key elements are defined as follows:
  • Let $\mathcal{D}_{S}=\{X_{S},P(X_{S})\}$ and $\mathcal{D}_{T}=\{X_{T},P(X_{T})\}$ be the source and target domains, respectively, where $X$ represents the input space and $P(X)$ the marginal probability distribution.
  • The tasks $\mathcal{T}_{S}=\{Y_{S},f_{S}(\cdot)\}$ and $\mathcal{T}_{T}=\{Y_{T},f_{T}(\cdot)\}$, with $Y$ being the output space and $f(\cdot)$ the target predictive function.
  • The goal is to improve $f_{T}$ by utilizing $f_{S}$, often requiring some labeled data in $\mathcal{D}_{T}$.
3.1.1 Fine-tuning
During fine-tuning, the model’s weights, initially learned from the source domain, are adjusted to better suit the target task (Figure 4(a)). This adjustment is done by continuing the training process on the target dataset, often with a smaller learning rate to make subtle changes to the weights. Fine-tuning in this context allows the model to adapt the knowledge acquired from the source domain to the specifics of the target domain and task, making it a quintessential example of inductive TL [50, 51, 52]. Figure 4(a) illustrates two methods of fine-tuning CNN: (i) to fine-tune the weights of the whole network and (ii) to fine-tune some certain layers. Figure 4(b) showcases a CNN instance of deep network adaptation, which modifies the distribution in the full connection layer by measuring the domain distance. For example in [53], the authors’ research highlights the significance of fine-tuning in TL for enhancing a pre-trained VGG19 and naïve inception model’s ability to detect kidney diseases from CT images. By adapting the DL model more specifically to this task, fine-tuning significantly boosts performance, achieving an impressive accuracy in identifying various kidney diseases. Similarly, the article [54] presents KidneyMultiNet, a web-based system for kidney disease detection using a hybrid ML model that combines two CNNs, DenseNet201 and Xception, and leverages TL techniques. This system, trained on a publicly available kidney CT scan dataset, achieved high performance and demonstrated the growing potential of advanced machine learning techniques in medical diagnostics. Furthermore, [55] introduces a methodology for kidney segmentation using TL with the U-Net architecture, employing a pre-trained EfficientNet as an encoder, which results in impressive segmentation precision and contributes to enhanced kidney tumor detection and treatment planning. Likewise, Kalluraya Puttur et al. [56] introduced a DL-based system for early KC detection using CT scans, building on the VGG16 architecture and enhanced with TL. This approach demonstrated the potential of DL in improving diagnostic accuracy and patient outcomes. In addition, the study [57] explores the transfer of a model trained on human MRI of ADPKD to animal models, effectively addressing challenges such as class imbalance and small datasets, highlighting the effectiveness of TL in cross-species medical imaging applications. Similarly, the study [58] introduces a TL model combining contrastive learning and ResNet-50 for renal pathology image classification, demonstrating how contrastive learning and TL can improve performance with minimal labeled data. In parallel, Sharma et al. [59] present an AI-based system using a modified InceptionV3 model for diagnosing kidney abnormalities, offering an autonomous and accurate diagnostic tool, particularly beneficial in areas with limited access to renal specialists. Additionally, Kaur et al. [60] propose a novel DL model based on EfficientNetB3 for classifying kidney diseases into four categories, demonstrating its high accuracy and practical utility in early diagnosis and treatment planning. In a similar vein, Prasher et al. [61] introduce a VGG16 TL model for diagnosing kidney stones using coronal CT images, highlighting its potential for clinical application by reducing dependence on manual diagnosis. Meanwhile, the study [62] investigates the role of activation functions in SqueezeNet’s performance for kidney disease classification, revealing that ReLU provides the best results, thereby emphasizing the influence of activation functions on medical image classification outcomes. Ghosh et al. [63] propose STREAMLINERS, a model that integrates fuzzy logic and TL-based feature extraction using DenseNet121 and ResNet101, achieving high accuracy on enhanced datasets. Similarly, Kumar et al. [64] develop a VGG16-based model that outperforms several alternatives in classifying kidney diseases, enhancing accuracy through fine-tuning and feature extraction, though it requires large labeled datasets. The study [65] proposes an image processing model for kidney stone classification using CT images, with image enhancement techniques like CLAHE and Laplacian filtering, and finds that ResNet101 performs best. This approach provides a reliable diagnostic tool for timely kidney stone detection. In the study [66], the authors introduce a smart model for kidney disease prediction, incorporating fine-tuning via a contrastive version of the wake-sleep algorithm, which effectively slows down kidney damage progression. Similarly, [67] presents a DL-driven segmentation framework that emphasizes the role of fine-tuning and novel loss functions to enhance model accuracy, particularly in kidney and renal tumor segmentation. Lastly, in [68], fine-tuning is used to adapt a 3D CNN (NephNet3D) to classify cell types in human kidney tissue, resulting in significant improvements in predictive performance.
Figure 4: Example of TL models used in KC: (a) Fine-tuning, and (b) DA.
3.1.2 Multi-Task learning (MTL)
Multi-task Learning aims at simultaneously learning multiple related tasks to improve generalization. It is mathematically described as:
  • Given a set of tasks $\{\mathcal{T}_{1},\mathcal{T}_{2},\ldots,\mathcal{T}_{N}\}$, the objective is to learn these tasks jointly by exploiting commonalities and differences.
  • This involves optimizing a shared representation $\Phi$ and task-specific functions
    $\{f_{T_{1}},f_{T_{2}},\ldots,f_{T_{N}}\}$.
  • The objective function is
    $\min_{\Phi,f_{T_{1}},\ldots,f_{T_{N}}}\sum_{i=1}^{N}\alpha_{i}\mathcal{L}_{T_{i}}(f_{T_{i}}(\Phi(X_{T_{i}})),Y_{T_{i}})$, where $\mathcal{L}_{T_{i}}$ is the loss function for task $i$.
The authors’ objective in [69] is to introduce an innovative semi-supervised multi-task learning method designed for forecasting the short-term progression of kidney disease using data from general practitioners’ electronic health records. They have shown that their methodology can: (i) track the progression of estimated glomerular filtration rate over time by enforcing a connection between sequential time periods, and (ii) harness valuable insights from patient data without labels, particularly when the data with labels are in limited supply.
In [70], the authors introduced a new multi-task learning approach aimed at assessing image quality, performing de-blurring, and classifying kidney diseases. The study analyzed immunofluorescence images from 1,608 patients, dividing them into 1,289 for training and 319 for testing. This innovative technique was not only effective in diagnosing four different kidney diseases using blurred immunofluorescence images, but it also demonstrated strong capabilities in two additional tasks: evaluating image quality and removing blur from images.
In [71], the authors introduce a multi-task 3D CNN designed for segmenting autosomal dominant polycystic kidney disease, which is identified by the continuous enlargement of renal cysts and stands as the most common severe hereditary kidney disorder. They successfully attained a high average dice score through this approach.
3.2 Transductive TL
The objective of transductive TL is to predict the labels $\{y_{j}^{T}\}_{j=1}^{n_{T}}$ for the instances in the target domain $\mathcal{D}_{T}$, leveraging the knowledge from the source domain $\mathcal{D}_{S}$, where:
  • A source domain $\mathcal{D}_{S}=\{(\mathbf{x}_{i}^{S},y_{i}^{S})\}_{i=1}^{n_{S}}$ with $n_{S}$ labeled instances, where $\mathbf{x}_{i}^{S}\in\mathbb{R}^{d}$ represents the feature vector of the $i$-th instance and $y_{i}^{S}\in\mathcal{Y}$ represents its label.
  • A target domain $\mathcal{D}_{T}=\{\mathbf{x}_{j}^{T}\}_{j=1}^{n_{T}}$ with $n_{T}$ unlabeled instances, where $\mathbf{x}_{j}^{T}\in\mathbb{R}^{d}$.
Transductive TL can be further categorized based on the strategies employed to bridge the gap between the source and target domains:
3.2.1 Domain adaptation
It focuses on adapting the source domain model to work effectively on the target domain by minimizing the domain shift. In other word, the aim of DA is to minimize the domain shift between the source domain $\mathcal{D}_{S}$ and the target domain $\mathcal{D}_{T}$ (Figure 4(b)) [72]. Techniques include:
  • Feature representation alignment: Aim to find a transformation $\phi$ that aligns the feature distributions of the source and target domains by minimizing a distance metric $d(\phi(\mathcal{D}_{S}),\phi(\mathcal{D}_{T}))$, such as the maximum mean discrepancy [73].
  • Domain-invariant feature learning: Learn a feature representation $\phi$ that is invariant to the change in domains, optimizing $\phi$ such that the predictive model $f(\phi(\cdot))$ achieves low error on both $\mathcal{D}_{S}$ and $\mathcal{D}_{T}$ [74].
  • Discrepancy minimization: Directly minimize the discrepancy between the source and target domain distributions using techniques like adversarial training, where a discriminator is trained to distinguish between the source and target domain features while the feature extractor $\phi$ is trained to fool the discriminator [75].
Instance weighting, feature transformation, and kernel methods are key techniques used in the DA process, aimed at addressing the issue of domain shift—where the data distribution in the source domain differs from that in the target domain [76]. These techniques facilitate the transfer of knowledge from a source domain to a target domain, helping to improve the performance of ML models on the target domain despite differences in data distribution [77].
  • Instance weighting: It assigns weights to instances in the source domain based on their relevance to the target domain. The goal is to give higher importance to source instances that are more similar to the target domain, thereby reducing the bias towards the source domain distribution [78]. Typically, given a source instance $(x_{i}^{S},y_{i}^{S})$, assign a weight $w_{i}$ based on its similarity to the target domain. The weighted loss for the source domain is then $\sum_{i=1}^{n_{S}}w_{i}\ell(f(x_{i}^{S}),y_{i}^{S})$, where $\ell$ is a loss function [79].
  • Feature transformation: It transforms the feature space to find a common subspace where the source and target domain data can be directly compared or combined. Methods include:
    • Subspace alignment: Find transformations $\phi_{S}$ and $\phi_{T}$ for the source and target domains, respectively, that align their subspaces, minimizing $\|\phi_{S}(\mathcal{D}_{S})-\phi_{T}(\mathcal{D}_{T})\|$ [80].
    • Manifold learning: Discover a low-dimensional manifold that both domains can be projected onto, learning a mapping $\phi$ that captures the intrinsic geometry of the data [81].
    • DL approaches: Use neural networks to learn a domain-agnostic feature representation, possibly with domain-adversarial components [82].
  • Kernel methods: They employ kernel functions to map the original feature space into a higher-dimensional space, where the source and target domains are more comparable. This approach facilitates finding a common representation that reduces the distributional differences between the domains [83]. Given a kernel function $k(\cdot,\cdot)$, transform instances $x_{i}^{S}$ and $x_{j}^{T}$ into a higher-dimensional space via $\phi$, where $\phi$ is implicitly defined by $k$. The goal is to learn a predictive model in this space that reduces the domain discrepancy [84].
These strategies are not mutually exclusive and can be combined to develop more sophisticated transductive TL models that effectively utilize the knowledge from the source domain to make accurate predictions in the target domain. The study [85] presents reciprocal domain adaptation network for KC diagnosis, which integrates TL with mutual model adaptation and domain model learning. While mutual model adaptation extracts rich features through global and local pyramid pooling, domain model learning enhances representation learning by combining semi-supervised, domain-independent features. Consequently, reciprocal domain adaptation network effectively addresses issues like data bias, domain shifts, and privacy, providing a promising solution for early KC detection. In a similar context, Canbay et al. [86] propose a privacy-preserving TL framework, incorporating differential privacy with CNNs. Although classification accuracy decreases with increased privacy settings, the framework ensures strong performance while protecting patient data, thus offering a secure solution for early kidney disease detection. Furthermore, the study by Kwon et al. [87] focuses on improving segmentation for autosomal dominant polycystic kidney disease using U-Net and TL with MRI data. By testing different training scenarios, the study demonstrates that single-shot fast spin-echo images yield the best results, reinforcing the utility of TL in medical segmentation. Similarly, Lopez et al. [88] present a two-step TL method using ResNet50 for classifying kidney stones in endoscopic and charge-coupled device images. This approach improves accuracy and outperforms traditional models. In [89], the authors propose a two-stage source-free unsupervised domain-adaptation framework for cross-modality abdominal multi-organ segmentations. The proposed method only requires a well-trained source model and an unlabeled target dataset, and does not need access to the source data. The experimental results highlight that this approach achieves satisfactory performance in adapting a labeled CT dataset to an unlabeled MRI dataset for abdominal organ segmentations, including the right and left kidneys. Additionally, in [90], the authors present a new DA approach called dimension-based disentangled dilated DA, designed to tackle the problem of domain shift for medical image segmentation tasks without the annotations of the target domain. By using adaptive instance normalization, the method encourages the content information to be stored in the spatial dimension and the style information in the channel dimension. They validate the proposed method for cross-modality medical image segmentation tasks on public datasets, including liver, right kidney, left kidney, and spleen.
3.2.2 Domain generalization
[…]
DG focuses on the challenge of developing models that can generalize well to unseen domains. In domain generalization (DG), a model is trained on data from multiple source domains to perform effectively on data from target domains that were not encountered during training. The core objective is to learn a domain-invariant representation that captures the underlying structure of the data, enabling the model to make accurate predictions across different, unseen domains [91]. In DG, the goal is to learn a model $f$ that generalizes well to any target domain $D_{T}$, given multiple source domains $\{D_{1},D_{2},\ldots,D_{K}\}$ [92]:
$\min_{f}\sum_{k=1}^{K}\mathcal{L}(f(X_{k}),Y_{k})+\lambda\Omega(f)$ (1)
In [93], the authors present an innovative approach utilizing neural style transfer to reduce model bias towards texture and intensity, aiming to prioritize shape characteristics. They conduct experiments using 200 T2-weighted MR images, half with fat-saturation and half without, to segment kidneys in patients with polycystic kidney disease. The results demonstrate a significant improvement in segmentation performance, indicating better alignment with actual kidney shapes.
In [94], the authors address the challenge of single-source DG in medical image segmentation, specifically noting the issue of domain shifts common in clinical image datasets. Traditional approaches, relying on global or random data augmentation, often produce samples that lack diversity and informativeness, inadequately representing the potential target domain distribution. To tackle this, the authors propose a novel data augmentation strategy for DG, inspired by the inherent class-level representation invariance and style mutability of medical images. They suggest that unseen target data could be represented as a linear combination of class-specific random variables, each adhering to a location-scale distribution. This concept is operationalized through the use of constrained bézier transformation for augmenting data at both global and class-specific levels, significantly enhancing augmentation diversity. Additionally, a saliency-balancing fusion mechanism is introduced to increase the informativeness of augmented samples by utilizing gradient information for orientation and magnitude guidance. The authors theoretically demonstrate that their augmentation strategy effectively limits the generalization risk for unseen target domains. Their saliency-balancing location-scale augmentation method notably outperforms existing state-of-the-art techniques in two single-source DG tasks, underscoring the potential of their approach to improve kidney and other medical image segmentation tasks by addressing the challenges posed by domain shifts.
3.2.3 Self-Taught learning
In self-taught learning, the model learns from unlabeled data that is not necessarily from the same distribution as the labeled data it was initially trained on. The key idea behind self-taught learning is to leverage large amounts of unlabeled data to improve the learning algorithm’s performance on a related task for which labeled data is scarce [95]. The process typically involves two main steps: first, extracting useful features or representations from the unlabeled data, and then, using these features to enhance the learning process on the labeled dataset. Self-taught learning enables the model to generalize better to new, unseen data by exploiting the underlying structure shared between the labeled and unlabeled datasets, even when they come from different domains. This approach is particularly valuable in scenarios where acquiring labeled data is expensive or labor-intensive, allowing for more efficient use of available data resources [96].
$\min_{f}\mathcal{L}(f(X_{S}),Y_{S})+\alpha\mathcal{L}_{AE}(X_{T})$ (2)
3.2.4 Few-Shot learning
[…]
FSL is a subset of transductive TL that focuses on the ability of a model to learn and make accurate predictions from a very limited number of training examples. The primary challenge in few-shot learning (FSL) is to design algorithms that can quickly adapt to new tasks or classes using only a few labeled samples, often referred to as "shots." This approach is crucial in scenarios where collecting large amounts of labeled data is impractical or impossible, such as in medical diagnosis, where rare diseases may only have a handful of known cases. FSL leverages prior knowledge, gained from related tasks or a broader dataset, to infer the new task’s structure with minimal data. Typically, In FSL, the model learns from a few examples in the target domain through a meta-learning approach:
$\min_{f}\sum_{\tau\in T}\mathcal{L}(f_{\theta+\Delta\theta_{\tau}}(X_{\tau}),Y_{\tau})$ (3)
This involves techniques such as meta-learning, where the model learns learning strategies from multiple tasks, or embedding learning, where data is transformed into a space where tasks with little data can be more easily learned. FSL aims to mimic human learning’s efficiency and flexibility, enabling models to generalize well from limited information.
In [97], the authors address the challenge of DA in DL for medical imaging, specifically in classifying kidney stones from endoscopic images taken under various conditions. They introduce a novel approach combining FSL with meta-learning. Initially, a self-supervised learning phase pre-trains the model to improve feature generalization across domains. Subsequently, meta-learning fine-tunes these features specifically for the domain of kidney stones, adapting the model to new, unseen domains effectively. This method significantly enhances the model’s generalization ability and demonstrates a promising direction for overcoming DA challenges in medical imaging applications.
In this study [98], the authors present a 3D few-shot segmentation framework tailored for precise segmentation of conditions such as KC, despite the scarcity of target organ annotation in training samples. A network resembling U-Net is developed to infer segmentation by understanding the relationship between 2D slices from support data and a query image, incorporating a bidirectional gated recurrent unit to ensure consistency in the encoded features among adjacent slices. Furthermore, they propose a TL approach to adjust the model to the specificities of the target image and organ by pre-testing updates using arbitrarily selected support and query samples. This model significantly outperforms existing few-shot segmentation models and rivals fully supervised models trained on more extensive target data.
Table 3: A summary of TL types for KC.
Ref Year Model-based TL TL type KC task Datasets Image features Metrics
[53] 2022 VGG19 and Naïve inception Inductive Fine-tuning Detection Private data 4000, CT ACC
[99] 2022 U-Net Inductive Fine-tuning Segmentation HuBMAP 15, CT Dice score
[67] 2024 U-Net Inductive Fine-tuning Segmentation Private data CT -
[68] 2023 3D-MS-RFCNN Inductive Fine-tuning Segmentation KiTS 1059, 512×512, CT ACC and dice score
[70] 2021 CNN Inductive MTL Classification Private data 1608, CT ACC and AUC
[71] 2018 3D CNN Inductive MTL Segmentation Private data 203, CT Dice score
[89] 2022 U-Net Transductive DA Segmentation Private data 30, CT and 20, MRI Dice score
[93] 2021 CNN Transductive DG Segmentation Private data 200, MRI Dice score
[94] 2023 U-Net with EfficientNet-b2 Transductive DG Segmentation Private data MRI Dice score
[97] 2022 DL Transductive FSL Classification Private data 600, CT ACC
[100] 2019 KNN Transductive FSL Prediction GDCD 654 samples ACC
[98] 2021 U-Net Transductive FSL Segmentation BCV 30, 3D, CT Dice score
[101] 2022 CNN Transductive FSL Segmentation Private data 3D, 20, MRI, 30, CT Dice score
4 Applications of TL in KC
TL has been applied in the field of KC diagnosis and prognosis using medical imaging and genomic data. Figure 5 illustrate the applications of TL in KC. Table. LABEL:table:2, presents a summary of TL research conducted in the field of kidney disease diagnosis. Some applications include:
Figure 5: The TL-based applications of for KC.
4.1 Cancer detection TL
TL is a method in DL that utilizes knowledge acquired from one task to enhance performance on a novel, but related, task, as outlined in [102]. Within the scope of analyzing medical images, TL enables the training of models on extensive datasets of analogous images, like radiology scans, before refining these models on a smaller set of medical images designated for a specific purpose, such as the detection of KC. To deploy TL for KC detection, the initial step involves gathering a comprehensive dataset of medical scans, such as CT or MRI images, which includes both healthy kidneys and those with cancerous growths. This collection serves as the foundation for pre-training a CNN to differentiate between normal and cancerous images. After the CNN has been pre-trained on the extensive dataset, it is then fine-tuned using a more focused collection of kidney images, precisely for KC detection. This fine-tuning adjusts the CNN’s pre-trained weights to refine its performance for this specific detection task. Throughout the fine-tuning phase, the CNN hones in on distinct characteristics within the kidney images that signal the presence of cancerous tumors, as illustrated in Figure 6. This process benefits from the insights gained from the larger dataset of similar images, allowing the CNN to learn how to identify KC in the smaller dataset with increased speed and accuracy [103].
Figure 6: An example that shows a workflow for classifying renal tumors using DL techniques. It starts with data augmentation, where images from public databases and internal data are processed through transformations like rotation, brightness adjustment, contrast enhancement, and scaling. These augmented images are then fed into various DL models, such as U-Net, ResNet, AlexNet, Inception-v3, and VGG-16. TL is applied to improve the model’s accuracy and reduce training time. The models then classify the images into categories: Adrenal Tumor or Normal Tissue. Additionally, multiomics data is integrated to enhance classification accuracy, incorporating both imaging and molecular data [104].
4.2 CT scan analysis, and segmentation TL
TL is a prominent strategy in DL, facilitating the adaptation of pre-trained models to novel tasks. This method has been particularly effective in the medical imaging domain for tasks such as the automatic segmentation applications (Figure 7) and classification of kidney tumors from CT scans. TL operates on the principle of leveraging the insights acquired from a model trained on a related issue trained on a similar dataset. This strategy is advantageous because it conserves time and resources that would otherwise be expended on initiating training from the ground up, while also enhancing accuracy through the transference of knowledge from the pre-trained model. Specifically, in the analysis of CT scans, TL has proven beneficial in refining pre-trained DL models for more precise and efficient identification and categorization of kidney tumors. Additionally, Gido et al. [105] present a DL approach for kidney and tumor segmentation using 3D CT data, demonstrating that fine-tuning is optimal for kidney segmentation, while mixed datasets are more effective for tumor segmentation. The authors in [99] emphasize the importance of hyperparameter fine-tuning, test time and train time data augmentations, and normalization in improving model performance for medical image segmentation tasks. The paper [106] introduces an ensemble TL-based model for kidney segmentation, incorporating multiple pre-trained CNNs and advanced data augmentation to enhance segmentation robustness and generalization. Similarly, Andreini et al. [107] adopt a cross-species TL strategy using MobileNet and SegNeXt for glomeruli segmentation, achieving strong performance even with limited annotations. Moreover, the study by Miao et al. [108] develops a CNN with TL (ResNet-18) for segmenting ultrasound images of renal tissue ablation, outshining conventional methods and performing comparably to manual segmentation by experts.
Figure 7: Example of the application of TL and segmentation in kidney cancer prognosis. It presents kidney segmentation on axial slices using a combined dataset model’s predictions in column 2, the model predictions without TL in column 3, and those with TL in column 4. Each row corresponds to a distinct test case, where yellow contours represent the ground truth kidney labels, and red contours indicate the model’s predictions [109].
Table 4: Summary of research conducted in the TL-based application of kidney disease.
Ref. Used method Dataset Methods description Best performance Limitation
[109] U-Net+TL Private data Proposed a fully automated 2D attention U-Net model for kidney segmentation on T1 weighted-nephrographic phase contrast enhanced DC:89.34$\pm$5.31 The model’s performance may still be limited by the relatively small size of the dataset
[110] CNN+CM +TL Private data Propose a cross-modal TL from CT to US to segment kidney organ based on CycleGAN, U-Net, U-Res, PSPNet, and DeepLab v3+ and CNN models. DC: 0.853 Limited training data for ultrasound Images
[111] CNN+TL Private data A 9-layer CNN based on the common U-Net. The architecture was developed and tested for the segmentation of non-sclerotic and sclerotic glomeruli of kidney. F1: 87.00
SEN: 93.00
Complexity of model.
[112] VGG16+TL Private data The U-Net was used to extract the kidney stone area. The VGG16 is used as the encoder of U-Net model to extract the semantic information. F1: 97.03 Pretrained features may limit model’s adaptation to endoscopic image characteristics.
[113] TL+CNN +BDRN CHOP Used a boundary distance regression network to learn kidney boundary ACC: 98.90 $\pm$0.6
DC: 0.94$\pm$0.03
Model struggles with US artifacts like noise, shadowing, and distortions.
[114] TL+SVM Private data support vector machine (SVM) classifiers and TL are used to classify diseased and normal kidneys. ACC: 87.00
SEN:86.00
Small dataset size affecting generalizability and robustness.
[115] DNN+TL GCDB Test DNN+TL model for classifying normal and abnormal categories of glomeruli ACC: 88.23
SEN: 88.23
Small dataset, TL models underperform, limited generalizability.
[116] TL+SVM CHOP Propose a TL-based method and SVM model to extract features from US kidney images to improve the CAKUT diagnosis in children. ACC:0.87$\pm$2.1
AUC:0.92$\pm$0.7
TL performance impacted by dataset imbalance and feature complexity.
[117] VGG16 Private data Development an intelligent system based on EANet, CCT, Swin transformers and Resnet, VGG16, and Inceptionv3 models to auto-diagnose kidney stones, cysts, and tumors. ACC: 99.30 Imbalanced dataset and limited modalities
[118] VGG16 with explainable AI (XAI) KUB X-ray images Developed a system for classifying kidney stones using TL and XAI to enhance transparency and decision-making. ACC: 97.41 Lacks fairness and explainability without XAI integration.
[56] VGG16 KiTS19 Dataset Predict KC from CT scans, integrating VGG16 with TL. ACC:97.5 Limited by dataset variability
[64] VGG16 Kaggle kidney disease TL for kidney disease classification. ACC:98.00 Requires large labeled data.
[61] VGG16 Coronal CT images of kidney stones Automated kidney stone diagnosis using TL model. ACC: 99.70 Performance depends on the quality and consistency of CT images.
[119] VGG18 Kidney ultrasound images Automated renal tumor classification using deep TL on ultrasound. SEN: 79.00
SPE:86.00
Limited by ultrasound image quality and operator skill dependency
[65] AlexNet, Xception, VGG19, ResNet101 CT images of kidney stones A TL-based model for kidney stone classification using CLAHE for enhancement, Laplacian filter for sharpening.. ACC: 98.15
Recall: 0.991
Limited to CT images, and may not generalize to other image modalities or conditions.
[86] Xception, ResNet50, InceptionResNetV2, VGG19 Priviat data Proposed a privacy-preserving TL framework for kidney disease detection using CNNs and differential privacy. ACC: 99.83
Recall 0.9983
Privacy-aware classification results degrade as noise increases
[108] ResNet-18 Ultrasound images CNN with TL segments ultrasound images of histotripsy.. DSC:85. Performance drops in early stages
[87] U-Net FIESTA and SSFSE MRI datasets TL applied for improved kidney segmentation using multi-sequence MRI data.. DC: 0.952 Limited to specific MRI sequences
[55] U-Net with EfficientNet as encoder KiTS19 Dataset Kidney segmentation using TL with U-Net architecture for CT images. DC:96.00
Jaccard index:94.40
Limited to kidney tumor segmentation
[105] nnU-Net KiTS19 and UTH datasets Explores kidney and tumor segmentation using mixed datasets and fine-tuning TL. DICE scores:0.964 Limited to kidney and renal tumor segmentation
[57] Modified UNet3 with residual layers MRI images of ADPKD kidneys TL applied to transfer a model trained on human ADPKD MRI images to rat and mouse ADPKD models.. DSC: 0.93$\pm$0.04 Class imbalance in target datasets
[120] Deep convolutional TL with SHAP Hunan Cancer Hospital and COIL-100 datasets Proposed an RCC prediction model incorporating SHAP for explainability and DCA for evaluating clinical utility. ACC:73.87 on RCC,
ACC:99.81 on COIL-100
Lower performance in real-world clinical datasets compared to public datasets.
[121] ACWGAN-GP, MobileNetV2 RenalUS images Developed an auxiliary diagnosis system for KC stage classification using DL. ACC: up to 90.1 Limited to four stages of KC.
[122] SSLSD-KTD with masked autoencoder KAUH-kidney, CT-kidney Combines self-distillation, self-supervised learning (SSL), and TL for better metrics with small datasets.. ACC:99.82 Limited by dataset size.
[54] CNN Kidney CT scan KidneyMultiNet combines DenseNet201 and Xception with TL for kidney disease detection. ACC: 99.92 Limited by the dataset
[85] reciprocal domain adaptation network (RDAN) Real-world KC datasets RDAN uses TL, MMA, and DML for accurate KC diagnosis. ACC: 96.94
Precision: 98.81.
Requires diverse datasets and may face challenges in clinical deployment across all environments.
[58] Contrastive learning(ResNet-50) TG-GATEs pathological image Proposed a TL model based on contrastive pretraining with mouse glomerulus images to improve human IgA nephritis image classification. ACC: 92.22
SEN: 92.74
Limited to glomerulus detection and classification
[123] Random Forest, Decision Tree, K-nearest Neighbor. Pediatric Ultrasound Kidney Images Introduces a BCOA to select optimal features from high-dimensional ultrasound images for KC detection. ACC:99.00
Precision: 100
The performance of the model may be highly dependent on the dataset size
[106] ResNet-50, VGG16, InceptionV3 KiTS19 Dataset Proposes an ensemble CNN model with TL for kidney segmentation. ACC:98
Precision: 0.91
Performance depends on image quality
[59] Fine-tuned InceptionV3 Priviat data Proposes AI system using TL(InceptionV3) for autonomous kidney disease diagnosis.. ACC: 96.52
Precision: 97.76
The model’s performance depends heavily on high-quality
[60] EfficientNetB3 Priviat data Proposes EfficientNetB3 TL model for automatic kidney disease classification. ACC:99.00 The model’s performance is influenced by the quality and diversity of the training dataset.
[62] SqueezeNet Kidney radiographs Analyzing activation functions’ impact on SqueezeNet classification. ACC: 51.17
SEN:90.55
Performance depends on activation function selection
[124] Xception Kidney images dataset AI-based system for accurate kidney disease prediction and classification.. ACC: 99.89 Requires large, well-labeled dataset for effective training.
[63] STREAMLINERS Kidney CT image dataset Fuzzy logic-enhanced CT images, ensemble learning, TL for tumor detection.. ACC: 99.25 Sensitive to noisy images and data imbalance.
[125] DarkNet19, InceptionV3, ResNet101 Kaggle kidney CT image Ensemble model using inductive transfer for kidney stone detection. ACC: 99.8 Limited dataset and performance degradation with noisy images
[88] ResNet50 with Two-step TL Endoscopic and CCD camera images Two-step TL for kidney stone classification. ACC:0.904$\pm$0.048 Small dataset and class imbalance
[107] MobileNet, SegNeXt Human Glomeruli, Mouse Glomeruli Cross-species TL for kidney glomeruli segmentation. Dice score:
90.43
Data differences between species
Note: The best performance are in (%).
4.3 Multi-modal TL-based KC
TL has been instrumental in enhancing the precision of KC diagnosis and prognosis by consolidating diverse data sources, including CT scans, MRI images, and gene expression data. This approach allows for a comprehensive analysis by leveraging the strengths of each data type. For instance, imaging data like CT and MRI scans offer visual insights into the physical state of the kidneys, identifying tumors’ size, shape, and location. On the other hand, gene expression data provides a molecular perspective, offering clues about the tumor’s aggressiveness and potential response to treatment. By applying TL, models pre-trained on large datasets from one domain (e.g., radiology imaging) can be fine-tuned with data from another domain (e.g., genomic data), facilitating a multi-faceted understanding of KC. This integration enables the models to learn from the complex interplay between the imaging characteristics of tumors and their molecular profiles, leading to more accurate and personalized diagnostic and prognostic assessments (Figure 8). The ability to draw upon pre-existing knowledge from various domains not only saves significant time and resources but also enhances the models’ predictive capabilities, making TL a powerful tool in the fight against KC.
Figure 8: An example of the use of multi modal data fusion in cancer prognosis, specifically for RCC. This approach integrates diverse data types such as radiological images, histology slides, and genomic information into a unified model to enhance diagnostic and prognostic accuracy. Panel (A) demonstrates survival analysis across different clinical stages (T, N, M stages, and grading), illustrating how integrating various data sources can provide deeper insights into disease progression. Panel (B) shows how radiological, morphological, and histological data are combined to analyze tumor characteristics, while Panel (C) outlines the fusion process, using imaging and genomic data to create predictive models that help in forecasting survival outcomes and guiding treatment strategies [126].
4.4 Explainable AI for TL-based KC
The potential of XAI in the diagnosis of KC presents a promising frontier in medical imaging and oncology [127]. Ahmed et al. [118] address this by applying a TL model with a pre-trained VGG16 architecture, enhanced with XAI techniques like layer-wise relevance propagation (LRP), achieving a remarkable accuracy in classifying KUB X-ray images as kidney stones or normal cases. This incorporation of XAI not only boosts the model’s performance but also its transparency, making the decision-making process more understandable to medical professionals. In the same direction, Han et al. (2021) tackle the prediction of RCC using deep convolutional TL and SHAP for model interpretability, achieving significant accuracies and providing valuable insights into model decisions [120]. This approach not only facilitates the understanding of how certain features affect the model’s prediction but also improves the clinical utility of the models by providing insights that are actionable in clinical practice.
5 Existing datasets and metrics
The efficacy of computer-aided detection (CAD) systems is significantly dependent on the dataset they are trained on, whether these systems employ TL techniques or other methodologies. The volume and integrity of the dataset not only affect its quality but also the performance of the resultant models. Yet, the acquisition of extensive medical imaging data presents challenges due to the intricate nature of imaging techniques and potential risks to patients. Moreover, the process of annotating such data demands the expertise of skilled radiologists to ensure precision, making high-quality, large-scale annotated datasets rare. In response to this limitation, research into novel data augmentation strategies, including photometric and geometric alterations, is underway to enrich existing datasets, thereby alleviating the issue of data scarcity. Enhancing the dataset’s quality via advanced imaging technique is another strategy to bolster the performance of models based on TL. These initiatives pave the way for enhanced model outcomes through the use of enriched data. There are several key datasets accessible to the public that offer insights into kidney tumors, namely: (i) The kidney tumor segmentation 19 (KiTS19) challenge dataset, offering CT scans for tumor segmentation algorithm assessment. (ii) The national cancer institute’s kidney renal clear cell carcinoma dataset, with gene expression and clinical details pertinent to CCRCC, the prevalent form of KC. (iii) The the cancer genome atlas (TCGA) kidney renal papillary cell carcinoma dataset, which provides genomic and clinical data for patients with PRCC. (iv) The dataset from radboud university medical center focusing on kidney tumors, featuring CT scans alongside segmentation masks for evaluating kidney tumor segmentation techniques. Each dataset is distinct in terms of the data it encompasses, the patient count, and the imaging techniques employed, as summarized in Table. LABEL:table:5, which centers on datasets related to kidney tumors.
Table 5: Summary of the most frequently used KC datasets.
Ref Available? Datasets Format RR Year Applications The statistics of dataset
[128] Yes [^1] https://archive.ics.uci.edu/ml/datasets/chronic_kidney_disease UCIMLR DICOM [129] 2018 Classify KC using SVM 400 cases
25 attributes
[130] 2019 Diagnosis of KC using ANN and SVM 400 patients
24 attributes
[131] 2020 Classify KC using MKSVM and FFOA 400 instances
25 attributes
[132] 2021 Predict KC using TL 300 patients
210 labels
[133] 2022 Kidney segmentation using RCNN
[134] Yes [^2] https://www.cancer.gov/tcga TCGA DICOM [135] 2018 Distinction between RCC using ANN and SVM 26 patients
[136] 2019 Prediction of mutation in RCC using ANN and RF 267 patients
[137] 2020 Differentiating high- from low-grade RCC using TL 390 patients
[138] 2021 Integrate CT and genomics features for RCC using SVM-RFE 205 patients
107 features
[139] 2022 Classify genomic characteristics in RCC using ML 575 cases
[140] Yes [^3] https://kits19.grand-challenge.org/ KiTS19 DICOM
NIfTI
[141] 2019 Segmentation of kidneys from CT images of kidney stone patients 153 patients
[23] 2020 Delimit the kidneys in CT images 210 patients
[132] 2021 Kidney segmentation 300 patients
210 labels
[142] 2022 Segment kidney and tumor regions 20 patients
210 images
[143] Yes [^4] https://www.ircad.fr/research/data-sets/ IRCAD DICOM
NIfTI
[144] 2019 Renal segmentation 363+128 patients
[145] Yes [^5] http://www.miccai.org/ MICCAI DICOM [146] 2019 Kidney segmentation 210 patients
Abbreviation: Relevant references (RR)
When discussing the advantages and limitations of the DL and TL frameworks, it is essential to specify the criteria used for their evaluation, including key performance metrics such as accuracy, precision, recall, F1 score, and computational efficiency (e.g., training and inference time, processing power requirements). These metrics provide a comprehensive assessment of the models’ predictive capabilities, balancing correctness (accuracy), sensitivity (recall), and the trade-off between precision and recall (F1 score). In the context of TL, selecting the right evaluation metrics is particularly crucial for understanding how well a pre-trained model adapts to a new domain and generalizes after transfer. This comparative analysis can highlight where the proposed frameworks excel and where improvements may be needed. Additionally, conducting comparative quantitative performance evaluations would be valuable, allowing direct comparisons with state-of-the-art methods and providing a clearer understanding of the frameworks’ relative strengths and weaknesses. Table LABEL:table:6 provides detailed explanations of commonly used evaluation metrics in TL, further aiding the assessment of their effectiveness in real-world applications.
Table 6: Detailed summary of evaluation metrics for TL
Metric Formula Detailed Insight
Accuracy $\frac{\text{TP}+\text{TN}}{\text{TP}+\text{FP}+\text{TN}+\text{FN}}$ Measures the proportion of correct predictions (both positives and negatives). Simple but not always reliable for imbalanced data.
Recall (True Positive Rate) $\frac{\text{TP}}{\text{TP}+\text{FN}}$ Measures how well the model identifies positive instances. Important when missing a positive instance (false negative) is critical.
Precision $\frac{\text{TP}}{\text{TP}+\text{FP}}$ Measures the accuracy of the positive predictions. Important when false positives (incorrect positives) are costly.
F1 Score $2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ Combines precision and recall into a single metric. Useful when there’s a need to balance both, especially with imbalanced datasets.
Specificity (True Negative Rate) $\frac{\text{TN}}{\text{TN}+\text{FP}}$ Measures the proportion of correctly identified negative instances. Important in scenarios where false positives are undesirable.
AUC - ROC Area under ROC curve Measures the model’s ability to distinguish between positive and negative classes. A higher AUC value indicates better discriminative ability across various thresholds.
6 Limitations and key challenges
TL can have several advantages in the context of KC, including securing annotated data in the healthcare sector is a significant challenge, often requiring considerable effort. TL plays a pivotal role in this context by facilitating the use of existing pre-trained models. These models, which have been educated on large and universally accessible datasets, can be adapted to develop specialized models for diagnosing specific conditions like KC. The advantages of applying pre-trained models are manifold. Firstly, these models, having been exposed to vast amounts of data, have the capability to significantly enhance performance on smaller, task-specific datasets, particularly in diagnosing KC. This approach not only leads to considerable savings in time and financial resources but also accelerates the training process. This is because fine-tuning a pre-trained model for a new task is much faster than starting training from scratch. Furthermore, TL reduces the need for extensive data and computational power for training models on specific tasks. In the realm of medical imaging, where missing or scarce data can pose significant challenges, especially for rare diseases like KC, TL offers a valuable solution by employing models that have been pre-trained on extensive datasets. This method effectively bridges the data gap, providing a way to address the issues of scarce data and domain shift that are prevalent due to the high variability in medical imaging data, making it especially beneficial for scenarios involving small datasets.
In addition there are several limitations to using TL for KC diagnosis:
The application of TL in diagnosing KC presents several challenges and considerations. One primary challenge is DA, which necessitates identifying a pre-trained model trained on a domain or task closely related to KC diagnosis. The effectiveness of TL hinges on the alignment between the original and target domains and tasks, with discrepancies potentially undermining the model’s ability to transfer knowledge effectively. Moreover, the risk of overfitting is heightened when fine-tuning pre-trained models on limited datasets, leading to models that perform poorly on new or unseen data. Fine-tuning entails careful adjustments of model layers, learning rates, and regularization methods to avoid such overfitting. Data scarcity for the new task can further complicate fine-tuning, potentially resulting in suboptimal performance. Additionally, TL can be computationally demanding, especially with large pre-trained models, posing scalability challenges. Interpretability issues also arise, as understanding the decision-making process of a model trained on a different task can be complex. The biases in the datasets used for training pre-trained models may carry over, affecting performance on certain data subsets. Moreover, these models, being trained on specific datasets, might not generalize well across different datasets, especially when data distributions vary significantly. Privacy concerns are another critical consideration, particularly when sensitive medical data must be shared with third parties for model training. The architecture of the pre-trained model also plays a crucial role in the success of TL ; an ill-suited architecture can hinder effective fine-tuning. Lastly, the substantial computational resources required for TL can make the use of pre-trained models costly, especially when employing multiple models or conducting several rounds of fine-tuning. These challenges underscore the complexity and nuanced considerations involved in applying TL for medical diagnosis tasks such as KC.
The studies reviewed in this article emphasize the effectiveness of TL for diagnosing KC, noting its computational efficiency and potential to outperform conventional ML and DL techniques. However, several key challenges still need to be addressed to enhance the performance of TL. This section aims to highlight the most significant unresolved issues and current concerns that are receiving considerable focus in the area of KC diagnosis:
6.1 Availability, adapting and adjusting models
Data accessibility plays a pivotal role in crafting and applying TL models for diagnosing KC. Nevertheless, several challenges complicate data accessibility in this arena. A primary issue is the scarcity of annotated medical imaging data specific to KC. This data, essential for TL model training, comprises images marked with details indicating the presence or absence of diseases like KC, yet it’s often limited in both amount and quality. Another obstacle arises from the inconsistencies in imaging techniques and data collection methods among various hospitals and healthcare facilities, which can introduce irregularities in the data for TL models, adversely affecting their performance and precision. Moreover, concerns related to privacy and security pose significant restrictions on medical imaging data’s availability for TL, given the sensitive nature of patient information embedded in such data, complicating its distribution among distinct entities. Additionally, data bias presents a critical challenge; TL models trained on skewed data might exacerbate existing inequalities in medical diagnoses and treatments. For instance, a model trained predominantly on data from male subjects might underperform with female patient images, reflecting and potentially worsening disparities.
The primary obstacle in TL involves identifying an appropriately pre-trained model that has been trained on a domain or task closely related to the new task at hand. It’s critical to evaluate how similar the original (source) and the intended (target) domains are, as well as the resemblance between the tasks involved [147, 148].
Adjusting a pre-trained model for a specific task, known as fine-tuning, presents its own set of challenges. It necessitates a strategic selection of which layers to adjust, alongside the determination of an optimal learning rate and the application of appropriate regularization strategies to prevent the model from becoming too narrowly adapted to the task [149].
6.2 Data scarcity
Data scarcity presents a significant challenge in the development and efficacy of KC diagnosis systems, impacting everything from image analysis to the training of DL models. Several studies have explored these challenges and proposed methods to mitigate the impact of limited data availability [150]. Peng et al. [151] explore how downsampling and compression affect renal image analysis. They found that data scarcity impacts diagnostic accuracy, highlighting the need for robust data handling techniques to ensure the integrity and usefulness of renal images in diagnosis. Similarly, Guo et al. [152] address the limited training data for kidney segmentation by using a cascaded CNN, demonstrating that innovative network architectures can partly overcome the challenges posed by sparse data.
In the context of clinical applications, Heller et al. [153] provide the kits19 dataset, which includes 300 kidney tumor cases with CT semantic segmentations and surgical outcomes, aiming to enhance research in KC by making rich datasets available. This effort underscores the critical role of accessible high-quality datasets in advancing early tumor detection and improving surgical outcomes. Flitcroft et al. [154] and Nassour et al. [155] both discuss the scarcity of reliable diagnostic biomarkers for KC, particularly in non-invasive strategies like urinary protein analysis and the evaluation of Lynch syndrome’s impact on KC. These studies highlight the gap in epidemiological data and the need for comprehensive datasets to develop effective diagnostic tools.
Technological advancements in handling data scarcity are also evident in DL applications. Alzu’bi et al. [156] and Bhattacharjee et al. [157] discuss the use of CNNs for kidney tumor classification and detection. They focus on optimizing DL approaches to perform effectively even with limited data, a crucial advancement given the rarity and variability of KC data. Lastly, Wang et al. [158] delve into the utilization of DL techniques for the imaging diagnosis of RCC. They emphasize how these advanced computational tools allow clinicians to diagnose and evaluate renal tumors more accurately and swiftly. One of the key strategies discussed is leveraging unlabeled data, which is a valuable solution for coping with data scarcity. By utilizing semi-supervised learning techniques, the potential of existing datasets can be maximized without the need for extensive labeled data.
6.3 Data bias and generalization
Data bias represents a significant obstacle in TL for KC diagnosis, impacting model accuracy and fairness. Bias can manifest in various forms, such as sampling bias, where training data may not accurately reflect the disease’s true distribution, potentially leading to population-specific biases [159]. Measurement bias skews data collection towards specific attributes, like age or gender, causing model predictions to favor these characteristics. Annotation bias arises from annotator prejudices, affecting the reliability of data labels [160]. To counteract these issues, TL models may utilize data augmentation to enrich training data variety and volume. Training models on data from diverse sources, like multiple hospitals, ensures a representative dataset of the disease’s distribution. Additionally, implementing fairness constraints and promoting algorithmic transparency can help mitigate bias, fostering the development of fair and unbiased models [161].
Generalization stands as a pivotal challenge in TL for diagnosing KC, denoting the model’s capacity to deliver accurate predictions on unseen data. Achieving robust generalization is essential, determining the model’s viability in practical scenarios [162]. A notable hindrance to strong generalization is the limited availability of annotated data. For many conditions, including KC, annotated data for training may be scarce, leading to a risk of overfitting, where the model overly specializes on training data, diminishing its performance on novel data. Additionally, data diversity issues, where training data may not accurately represent the disease’s distribution, can impair generalization [163]. For instance, data sourced from a singular hospital might not mirror the broader disease prevalence across different locales. To mitigate these issues, TL models for KC diagnosis may incorporate data augmentation strategies to broaden the training data’s size and variety. Training on data from varied sources, like multiple healthcare facilities, ensures a more representative dataset. Furthermore, employing regularization methods like dropout aids in curbing overfitting, enhancing the model’s generalization capabilities [164].
6.4 Privacy concerns, expertise shortage and accuracy instability
Privacy concerns emerge prominently in developing and applying TL models for KC diagnosis due to the necessity of accessing sensitive medical information, triggering significant privacy and security considerations [159]. A key challenge is safeguarding medical data privacy and security, demanding stringent measures against unauthorized access and potential breaches. Managing medical data with utmost care to align with privacy regulations is crucial [34]. The annotation process also presents privacy challenges, necessitating judicious annotator selection and bias minimization efforts. Employing privacy-preserving methods, such as differential privacy, is essential to protect patient confidentiality during annotation. Additionally, securing TL models against unauthorized interventions and ensuring their resilience is paramount. Providing patients with transparent data usage information and granting them control over their data usage decisions is vital for maintaining trust and privacy integrity [165].
The creation and implementation of TL models for KC diagnosis necessitate profound expertise in medical imaging and ML, yet expertise scarcity poses several challenges [166]. The limited pool of experts hampers the resources and knowledge necessary for these models’ development and application. Interdisciplinary collaboration, essential between medical and ML experts, is often obstructed by differing methodologies and communication barriers [167]. Adhering to medical imaging and mML best practices and standards is imperative for the responsible development and deployment of TL models. Ensuring patient safety and efficacy of these models requires a deep understanding of patient needs and the development of models that positively impact patient outcomes.
Improving TL model performance often involves utilizing pre-trained models on extensive datasets, such as those trained on the ImageNet database. Nonetheless, limited data availability can lead to accuracy instability. Various regularization techniques, such as SPAR, have been explored to stabilize accuracy by moderating the target network’s outer layer weights with reference to a starting point [168]. An innovative regularization method, DL transfer with attention, was introduced, focusing on preserving the source model’s outer layer outputs rather than merely constraining the model’s weights. DL transfer with attention minimizes empirical loss and aligns the two models’ outer layer outputs by selectively constraining feature maps identified by a supervised learning-trained attention mechanism .
6.5 Knowledge gain and data annotation
Evaluating the knowledge gained by the TL model and its performance outcomes is crucial. Consequently, considerable research has been directed towards establishing precise metrics for assessing the effectiveness of these models and the extent of knowledge they acquire, aiding in the elucidation of complex issues [161]. Proposed metrics include transmission error, transmission loss, transmission ratio, and in-range ratio. Nevertheless, the specificity of these metrics often falls short of expectations, highlighted by the ongoing uncertainty regarding outcomes when other TL-based models are employed, such as in KC detection scenarios, where model performance might be suboptimal.
Data annotation plays an essential role in developing TL models for diagnosing KC. Yet, this process faces several significant hurdles. A primary issue is the considerable cost and time required to annotate extensive datasets accurately. The process demands substantial expert input and resources, making it both time-intensive and costly. Another obstacle is the absence of standardized practices in data annotation, complicating the comparison and amalgamation of datasets from diverse origins and potentially diminishing the TL models’ impact. Ensuring the annotations’ quality and precision further necessitates meticulous validation and quality control, adding to the resource and time expenditure. Moreover, achieving representativeness in annotations, ensuring they accurately reflect the target demographic, involves careful annotator selection and concerted efforts to minimize bias throughout the annotation process.
6.6 Negative transfer and overfitting
Negative transfer in the context of learning about KC refers to the difficulties that individuals may experience when acquiring new knowledge or skills related to the condition and its treatment [169, 170]. (1) Overgeneralization: Individuals may apply learned information too broadly, which may result in an incorrect or incomplete understanding of the condition and its treatment. (2) Overreliance on prior knowledge: Individuals may rely too heavily on prior knowledge about cancer or other medical conditions, which may not be applicable to KC.(3) Incompatibility between old and new learning: Prior learning about cancer or other medical conditions may be incompatible with new information about KC, leading to confusion and hindering the acquisition of new knowledge. (4) Resistance to change: Individuals may be resistant to changing their existing beliefs and attitudes about cancer, leading to difficulty in adapting to new information about KC. (5) Misapplication of prior knowledge: Individuals may misapply prior knowledge, leading to an incorrect or inefficient understanding of the condition and its treatment.
Overfitting, especially in the domain of KC diagnosis, describes the situation where a predictive model becomes excessively tailored to the training dataset, undermining its capability to perform accurately on new, unseen data. This often results from training a model with an overly complex set of parameters, rendering it over-specialized and ineffective for general prediction or diagnosis purposes [171, 172]. In the realm of KC, overfitting may result in inaccurate diagnoses, an increase in false positives, or diminished disease detection sensitivity, posing significant risks to patient management and therapy. Implementing measures such as regularization to curb model complexity and employing cross-validation to evaluate model efficacy on novel data are critical strategies to prevent overfitting.
The study by [173] introduces innovative regularization techniques, including large margin loss, focal loss, adversarial training, mixup, and data augmentation. These methods aim to correct the logit shift in underrepresented classes and address class imbalance by scrutinizing network behavior, providing a robust approach to mitigate overfitting.
Furthermore, [174] details the use of Five-fold cross-validation as a means to assess the performance of AI algorithms, specifically in the segmentation process involving seed placement and bounding box techniques for lesion patch extraction, coupled with deep CNN for distinguishing clear CCRCC from renal oncocytoma.
7 Future research directions
Despite achieving high accuracy, prior studies highlight the need for improved determination of KC. Researchers suggest employing less conventional algorithms to enhance early-stage detection. Enhancing TL-based models for robust, unbiased KC identification across stages is crucial. Integrating feature selection with DL models and expanding datasets from various sources may support this goal. Further refinement of algorithms for related classification tasks is also needed. Potential future directions for TL in KC diagnosis are outlined in the following subsections.
7.1 Blockchain technology
In the context of KC diagnosis, TL plays a vital role in enhancing the effectiveness of blockchain technology. TL allows DL models to be trained on large, diverse datasets of medical images, such as CT scans or MRIs, for detecting tumors. Initially, the model is trained on a broad range of cancer data, and later refined using a more specialized dataset focused on KC. This process helps the model to become highly accurate in detecting KC-specific characteristics. Blockchain technology complements TL by providing a secure, transparent, and immutable platform to store and share sensitive medical data. It ensures that patient records are not only protected but also verifiable, offering patients control over their data and granting permissions to authorized medical professionals for analysis. Through blockchain’s decentralized nature, researchers and healthcare institutions can securely share and collaborate on medical data, enabling the use of TL-based models across multiple centers. This integration enhances diagnostic accuracy, reduces bias, and improves patient care outcomes by allowing for more precise, timely, and collaborative diagnoses of kidney cancer [175, 34].
7.2 Adversarial training
The burgeoning field of adversarial training offers a compelling avenue for enhancing the diagnosis, prediction, and segmentation of KC. This approach leverages the strengths of generative adversarial networks (GANs) and adversarial training methods to address some of the inherent challenges in medical imaging, such as data scarcity, variability in tumor appearance, and the robustness of diagnostic models against adversarial examples. Furthermore, TL plays a pivotal role in overcoming the challenge of limited labeled data in medical imaging by enabling the use of pre-trained models on large-scale datasets, which can be fine-tuned for specific tasks like KC detection. TL allows models to capture essential features from general medical imaging datasets and adapt them to the nuances of KC, thus improving diagnostic performance and model generalization. This integration of adversarial training with TL has the potential to significantly enhance the accuracy and robustness of KC diagnosis models, especially in clinical environments with limited annotated datasets [176, 177, 178, 179].
7.3 Compression and denoising based on TL for KC
A future direction for improving KC diagnosis and treatment could involve further advancements in TL. These innovations could enhance both medical image compression and tumor detection processes. Specifically, refining TL techniques for compressing KC images could ensure that critical diagnostic information is preserved while reducing computational and storage demands. Future research could explore more advanced model compression strategies, dynamic loss functions, and real-time data sharing to support collaborative decision-making among healthcare professionals. By combining these strategies, the goal would be to create faster, more accurate, and scalable solutions for detecting and diagnosing KC, leading to better patient care and treatment outcomes [180, 181, 182, 183]. Moreover, employing efficient pretrained denoising DL techniques—such as the one presented in [184]—can significantly improve the quality of medical images prior to analysis, thereby enhancing the accuracy and reliability of KC detection and diagnosis.
7.4 Quantum computing
Quantum computing represents a nascent and burgeoning domain poised to transform our approach to computational challenges, including those within ML, with the potential to revolutionize various aspects of knowledge discovery. Despite being at an early developmental stage, quantum computing harbors the promise of significantly enhancing the efficacy of TL models in the diagnosis of diseases like KC [185]. Employing quantum computing in TL for KC diagnosis entails harnessing quantum algorithms to tackle tasks that would otherwise be impractical with classical computers, notably optimization and linear algebra computations. These algorithms offer avenues to bolster TL model performance by curtailing computational overheads and enhancing model accuracy. Additionally, quantum computing holds promise in mitigating classical computing constraints encountered in medical imaging, such as the demand for substantial computational resources and memory for processing voluminous datasets. By leveraging quantum computing, the duration required for training and inference can be slashed, fostering swifter diagnoses and improved patient outcomes [186, 187, 188].
7.5 DA for clinical data
[…]
7.6 UL, and SSL-based TL for KC
[…]
UL, SSL, and TL for KC diagnosis can provide substantial benefits by enhancing diagnostic accuracy with limited labeled data. unsupervised learning (UL) can uncover hidden patterns and clusters in medical images, identifying previously unnoticed features that distinguish cancerous tissues. SSL can utilize large amounts of unlabeled data by generating pseudo-labels through pretext tasks, reducing the need for manually annotated datasets. TL further amplifies these techniques by enabling models trained on one dataset to be fine-tuned and adapted for KC detection in different clinical settings or patient populations, ensuring better generalization and scalability. Together, these methods can significantly improve early detection, reduce the reliance on expert annotation, and enable the development of more robust, efficient, and scalable AI systems for KC diagnosis [122, 190].
7.7 Federated TL for KC
A promising future research direction involves advancing the integration of TL and federated learning (FL) to tackle the complex issue of identifying kidney abnormalities in medical imaging. This approach could enhance diagnostic capabilities within decentralized healthcare systems while maintaining patient data privacy. Future studies may focus on refining the federated TL architecture to enable collaborative learning across geographically dispersed renal imaging datasets, allowing healthcare providers to share insights without compromising data ownership or confidentiality. Further exploration of this methodology could lead to improved precision in renal abnormality detection by adapting TL models to better capture the unique characteristics of renal imaging data. Additionally, continued research could enhance the efficiency and security of this system, facilitating more effective, privacy-preserving, and decentralized healthcare solutions. By building on this convergence of deep learning, federated learning, and TL, future work could significantly transform the landscape of medical image analysis, fostering a new era of collaborative, secure, and adaptive healthcare [191, 192].
7.8 Self-paced learning for KC
TL has proven to be a powerful approach for improving medical diagnostic systems, especially in the complex field of KC diagnosis. TL leverages knowledge gained from pre-trained models on large datasets to enhance the performance of models on specific tasks, such as KC detection. By incorporating a strategy that focuses on progressively more complex or informative cases as the model improves, TL adapts to the nuances of the data, helping to address challenges like data variability and imbalance. This adaptive learning process can greatly benefit KC diagnosis, allowing the model to effectively handle the complexities of medical imaging and improve diagnostic accuracy over time [193].
7.9 Tranformers and Large language models for KC
The integration of transformer architectures, such as bidirectional encoder representations from Transformer (BERT) and connectionist temporal classification (CTC) among others [194, 195], into the realm of KC diagnosis and segmentation represents a cutting-edge approach that holds significant promise for advancing medical imaging analysis. Drawing insights from several studies, such as those referenced, the potential applications of transformers in this domain are evident. One of the primary advantages of transformers lies in their ability to capture long-range dependencies within the data, making them particularly suitable for tasks involving complex relationships and patterns, as often encountered in medical imaging. The referenced studies showcase various transformer-based models tailored specifically for KC diagnosis and segmentation.
For instance, the utilization of convolution-and-transformer networks (COTRNet) proposed by Shen et al. [196] offers an end-to-end solution for kidney tumor segmentation, demonstrating the effectiveness of combining convolutional and transformer-based architectures in this context. Similarly, Islam et al. [117] leverage vision transformers for the automatic detection of kidney cysts, stones, and tumors from CT radiography, highlighting the versatility of transformer models across different types of kidney abnormalities.
Furthermore, the study by Wang et al. [197] showcases the synergy between transformers and CNNs for renal parenchymal tumors diagnosis. By stacking transformers and CNNs, the model efficiently captures both long-range dependencies and low-level spatial details, enhancing its performance in identifying and diagnosing kidney lesions. Moreover, transformer-based methods, such as TA-UNet3+ proposed by Hu [198], introduce innovative attention mechanisms and encoder-decoder architectures to improve kidney tumor segmentation accuracy. Similarly, Qian et al. [199] present a hybrid network combining nnU-Net and Swin Transformer, emphasizing the critical role of precise delineation and localization in KC diagnosis and treatment planning. Additionally, transformers exhibit promise in histopathological image analysis for subtype classification of RCC. Gao et al. [200] demonstrate the effectiveness of instance-based vision transformers in subtyping papillary RCC, providing valuable insights for histopathological diagnosis and personalized treatment strategies.
[…]
LLM and generative chatbots hold significant potential for transforming KC diagnosis in several ways [201, 202]. By integrating vast medical libraries and research data, large language model (LLM) can provide healthcare professionals with up-to-date information, diagnostic suggestions, and treatment options based on the latest studies and clinical guidelines [203, 204]. Moving on, generative chatbots can interpret patients’ descriptions of their symptoms in natural language, aiding in the early detection of KC by flagging potential warning signs and suggesting further diagnostic tests [205, 206, 207]. Typically, these models can deliver personalized information to patients about KC, including explanations of their condition, possible treatment pathways, and what to expect during each treatment option, improving patient understanding and engagement in their care plan [208, 209, 210].
Besides, LLM can analyze vast amounts of data from EHRs to identify patterns or risk factors associated with KC, potentially uncovering new insights into its diagnosis and progression [211, 212]. With training, LLM can assist radiologists by providing a second opinion on imaging studies, such as CT scans and MRIs, identifying areas that may warrant a closer look for signs of KC [213, 214, 215]. Additionally, chatbots can handle routine inquiries and patient triaging, allowing medical staff to focus on more complex tasks and patient care, thereby increasing the efficiency of clinical workflows [216, 217]. LLM can integrate and analyze information from various sources, including genomic data, radiographic images, and clinical notes, offering a holistic view of a patient’s health status and aiding in the accurate diagnosis of KC [218, 219]. Moreover, generative chatbots can facilitate remote monitoring of patient’s symptoms and treatment responses, providing timely advice and enabling telehealth consultations with specialists [220, 221].
8 Conclusion
The integration of TL and DA techniques marks a pivotal advancement in enhancing the accuracy, efficiency, and robustness of KC diagnosis. TL plays a critical role in overcoming the limitations of traditional DL approaches by enabling models pre-trained on large, diverse datasets to adapt effectively to smaller, specialized datasets, where annotated data is often scarce—a persistent challenge in medical diagnostics. This knowledge transfer not only reduces the computational burden of training complex models from scratch but also significantly improves diagnostic precision, even when fine-tuning with limited kidney-specific data. Complementing this, DA techniques address the critical issue of data distribution shifts, ensuring that diagnostic models remain reliable and accurate amidst the dynamic and evolving nature of medical imaging data. Through a systematic and comprehensive analysis of TL-based frameworks for KC detection, this study provides a rigorous evaluation of key contributions, identifies pressing challenges, and proposes actionable solutions to push the boundaries of the field. It underscores the transformative potential of TL in tackling major hurdles, including the computational complexity of DL, the scarcity of annotated medical data, and the variability between training and testing datasets, thereby enhancing diagnostic reliability and precision. Despite these advancements, significant challenges persist. Many DA models rely on the assumption of homogeneity in feature distributions between source and target domains—an assumption that often fails to hold in real-world medical scenarios, where data heterogeneity is the norm. Recent research has begun to explore heterogeneous DA approaches, which relax these assumptions and open new avenues for innovation. Looking ahead, future research should prioritize the optimization of these heterogeneous DA methods to improve their applicability in KC detection, fostering the development of more robust, generalizable, and scalable diagnostic systems. Furthermore, efforts should focus on integrating these techniques into clinical workflows, ensuring seamless translation from research to practice. By addressing these challenges, the synergy of TL and DA has the potential to revolutionize precision medicine, driving transformative changes in clinical practice, improving patient outcomes, and advancing the frontiers of oncology research. However, some limitations of this review must be acknowledged. The studies reviewed primarily focus on a narrow subset of methods, often omitting emerging techniques in the broader TL and DA fields that may offer significant improvements. Additionally, there is a heavy reliance on studies with limited clinical validation, which could impact the generalizability of the findings to real-world medical environments. Some studies also show methodological inconsistencies in how TL and DA techniques are applied, making direct comparisons challenging. Furthermore, the review has been constrained by the availability of high-quality, peer-reviewed papers, limiting the inclusion of non-published, yet potentially valuable, research. Future reviews should strive to include a broader range of studies, especially those that explore new, diverse, and interdisciplinary approaches to TL and DA in medical imaging.
Declarations
  • The authors declare no conflicts of interest.
  • Funding: The authors declare that there is no fund for this project
  • Data availability: There is no data available for this project
  • Ethics approval and consent to participate: Not Applicable.